### v1

# Quantum Simulation of Spin-Boson Models with Structured Bath

Abstract
The spin-boson model, involving a spin interacting with a bath of quantum harmonic oscillators, is a widely used representation of open quantum systems. Trapped ions present a natural platform for the quantum simulation of such models; however, previous experiments have been limited to simulating a scenario where a spin is coherently coupled to bosonic modes, which neglects capturing the dissipation of the bath. In our work, we perform quantum simulations of spin-boson models with structured baths using the motional states of trapped ions. We demonstrate the capability for adjusting the bath’s temperature and continuous spectral density by adding randomness to the control parameters. Subsequently, we simulate the dynamics of various spin-boson models with spectral densities composed of up to three Lorentzian peaks. The experimental outcomes closely align with theoretical predictions, suggesting the advantage of adding the trapped-ion system’s experimental noise for simulating open quantum systems.
Every quantum system inevitably interacts with the surrounding environment. Such interactions can significantly influence the dynamics of the system. A paradigmatic model of non-Markovian open quantum systems that captures the quantum nature of both the system and the environment is the spin-boson model Leggett et al. (1987), which is ubiquitous in the study of condensed-matter physics Caldeira and Leggett (1983); Weiss et al. (1987), chemical reactions Garg et al. (1985); Gilmore and McKenzie (2005), and biological light-harvesting complexes Huelga and Plenio (2013); Jang and Mennucci (2018). Experimentally realizing the dynamics of spin-boson models in a programmable fashion is an interesting challenge for quantum simulators Mostame et al. (2012); Leppäkangas et al. (2018); Magazzù et al. (2018); Kim et al. (2022).
Trapped ions provide a highly controllable system of spins and motional modes, and thus are a natural platform for simulating spin-boson models. The bosonic bath can be simulated using either a large number ($\gtrsim 50$) of motional modes Porras et al. (2008) or a few dissipative motional modes Lemmer et al. (2018); Schlawin et al. (2021); MacDonell et al. (2021); Kang et al. (2024). While significant experimental progress has been made in controlling the qubits and motional modes together Jia et al. (2022); Chen et al. (2023), past works in this direction are limited to simulating a spin either coherently coupled to up to two bosonic modes Gorman et al. (2018); MacDonell et al. (2023); Whitlow et al. (2023); Valahu et al. (2023). These previous studies did not consider the effects of dissipation, which is often described by the coupling of a system to a continuum of bath modes. Dissipation of the environment determines the system’s long-term behavior, such as reaching a steady state, and therefore is essential to incorporate for accurately modeling open quantum systems. Note that a recent study Wang et al. (2024) simulated a spin dissipatively coupled to many modes albeit with limited tunability.
In this work, we perform quantum simulations of spin-boson models with continuous spectral densities using trapped ions and their motional modes. Specifically, we simulate spectral densities composed of several Lorentzian peaks, which is characteristic of baths in light-harvesting protein complexes Olbrich et al. (2011); Chandrasekaran et al. (2015); Lee and Coker (2016); Kim et al. (2018); Tamascelli et al. (2019). Furthermore, we simulate the sub-Ohmic, Ohmic, and super-Ohmic spin-boson models Leggett et al. (1987), which are ubiquitous in various open quantum systems ranging from nanomechanical devices Seoanez et al. (2007); Galland et al. (2008); Chen et al. (2011) and superconducting circuits Magazzù et al. (2018); Thorwart et al. (2004); Kaur et al. (2021); Gulácsi and Burkard (2023) to molecules in which proton-transfer reactions occur Cukier and Morillo (1989); Ohta et al. (2006); Craig et al. (2007). Finally, we simulate the vibrationally-assisted energy transfer (VAET) Gorman et al. (2018) in a simple molecular model where dephasing is added to the vibrational mode. The various spin-state population dynamics, which display behaviors such as damped oscillations, coherent revivals, and relaxation to a steady state, are accurately captured by our quantum-simulation results.
Capturing the dissipative behavior of the spin-boson model, as well as tuning the bath’s temperature, is enabled by adding randomness to the control parameters such as laser frequency and phase. Remarkably, adding random frequency offset to the control of trapped ions’ motional modes is directly analogous to the leading source of noise of our experimental setup, namely, motional dephasing Kang et al. (2023a). Thus, motional dephasing noise of the hardware can essentially be used by our quantum simulations, providing accurate results even when the experimental time for simulating the time evolution is comparable to the coherence time of the motional modes. While the quantum simulations presented here are limited to the weak spin-bath coupling regime, our demonstration of adding or using noise Guimarães et al. (2023); Leppäkangas et al. (2023) in analog quantum simulations is an important ingredient for potential quantum advantage in simulating open quantum system models that are intractable using classical simulation methods MacDonell et al. (2021); Kang et al. (2024).
Results
Models
The spin-boson model is described by a spin coupled to a continuous bath of quantum harmonic oscillators. The Hamiltonian is given by ($\hbar$ is set as $1$ for simplicity)
[…]
A spin-boson model with structured bath is often characterized by a spectral density $J(\omega)$ composed of multiple Lorentzian peaks, each of which can be represented by a noisy harmonic oscillator Lemmer et al. (2018). Specifically, we consider a spin coupled to dephased oscillators (hereafter referred to as the dephased spin-oscillator model). The single-oscillator case is described by the Hamiltonian
$\hat{H}_{SO}=\hat{H}_{S}+\frac{\kappa}{2}\hat{\sigma}_{Z}\otimes(\hat{b}+\hat{b}^{\dagger})+\nu\hat{b}^{\dagger}\hat{b}$ (2)
and a single Lindblad operator $\hat{L}_{1}=\sqrt{\Gamma}\hat{b}^{\dagger}\hat{b}$, where $\hat{b}$ is the annihilation operator of the oscillator, such that the composite state $\rho$ of spin and oscillators follows the Lindblad master equation $\dot{\rho}=-i[\hat{H}_{SO},\rho]+\sum_{j}(\hat{L}_{j}\rho\hat{L}_{j}^{\dagger}-\frac{1}{2}\{\hat{L}_{j}^{\dagger}\hat{L}_{j},\rho\})$ Lindblad (1976). In a reasonable regime of parameters ($\Gamma<\nu/2$, $\beta\ll 2\pi(\Gamma/2)^{-1}$) Lemmer et al. (2018), this bath of dephased oscillators is assigned a spectral density composed of Lorentzian peaks sup
$J_{\rm Lo}(\omega)=\kappa^{2}\Big{(}\frac{\Gamma/2}{(\Gamma/2)^{2}+(\omega-\nu)^{2}}-\frac{\Gamma/2}{(\Gamma/2)^{2}+(\omega+\nu)^{2}}\Big{)},$ (3)
where we assume that initially the spin is in the |$\ket{0}$ state and the oscillator is in the thermal state of average phonon number $\bar{n}(\nu)$.
The spectral density of the bath determines the dynamics of the spin up to leading order in $\kappa T$, where $T$ is the evolution time Breuer and Petruccione (2002); sup . Therefore, the spin dynamics of the dephased spin-oscillator model approximately match those of the spin-boson model with $J(\omega)=J_{\rm Lo}(\omega)$ in the weak-$\kappa$ regime, which straightforwardly extends to the model with multiple oscillators. In particular, for the unbiased ($\epsilon=0$) spin-boson model, which is of particular interest Leggett et al. (1987); Wang and Thoss (2008); Duan et al. (2017); Strathearn et al. (2018), the equilibrium donor-state (|$\ket{0}$) population is $1/2$ for both dephased spin-oscillator models and spin-boson models. Thus, for weak spin-bath coupling, the population dynamics of the two models show an excellent match at all times; the match in short (long) time scale is achieved due to matching spectral densities (equilibrium populations). Unless specified otherwise, we use $\epsilon=0$; results for $\epsilon\neq 0$ can be found in Fig. 5.
A more commonly considered model is a spin coupled to damped oscillators (hereafter referred to as the damped spin-oscillator model); for example, the single-oscillator case is described by the Hamiltonian $\hat{H}_{SO}$ and a pair of Lindblad operators $\hat{L}_{1}=\sqrt{\Gamma(\bar{n}(\nu)+1)}\hat{b}$ and $\hat{L}_{2}=\sqrt{\Gamma\bar{n}(\nu)}\hat{b}^{\dagger}$ Lemmer et al. (2018); Schlawin et al. (2021). For this model, a stronger result holds: the spin dynamics are non-perturbatively equivalent to those of the spin-boson model with the same spectral density, up to all orders in $\kappa T$ Tamascelli et al. (2018); sup . However, realizing damped oscillators on trapped ions requires sympathetic cooling on a chain of multiple species of ions Lemmer et al. (2018); Schlawin et al. (2021), which is challenging to perform together with coherent operations that simulate the time evolution. Thus, we instead use dephased oscillators to represent the structured bath, as dephasing can be readily performed using randomized operations on trapped ions as described below.
Experimental results
The dephased spin-oscillator model can be simulated using trapped ions by applying spin-dependent kick (SDK) operations, which induce the coupling between the trapped-ion qubit and the motional mode (see Methods for details). Also, by adding randomness to the control parameters of the SDK operations, we can (i) prepare the thermal state and (ii) simulate the dephasing of the motional mode, by averaging the results over many random trials. The key idea is that noisy (heated or dephased) evolution, described by the Lindblad master equation, is equivalent to an average of many coherent evolutions, each with respect to a stochastic Hamiltonian Cai and Barthel (2013); Chenu et al. (2017).
Figure 1: Demonstration of the tunability of the bath’s temperature and spectral linewidth. (a) Dynamics of the coherent spin-oscillator model $\hat{H}_{SO}$ with various values of average phonon number $\bar{n}$ of the oscillator’s initial state. (b) Dynamics of the spin-boson model with $\bar{n}=0$, $J(\omega)=J_{\rm Lo}(\omega)$ and various values of $\Gamma$. For both (a) and (b), $\epsilon=0$, $\nu=\Delta$, and $\kappa=0.1\Delta$. Lines (dots) represent theoretical predictions (experimental data) and error bars denote the standard deviation over 20 trials, each trial performed with randomly drawn set of parameters and repeated 100 times. (c) Revival amplitude versus $\bar{n}$. Predictions of the original (solid line) model and the noise-aware (dashed line) model are compared with the experimental results (circles). (d) Fitted values of $\bar{n}$, where the experimental data are fitted to the predictions of the noise-aware model. Solid line depicts $y=x+\bar{n}_{0}$, where $\bar{n}_{0}=0.036$ is the expected offset in the phonon number due to imperfect cooling. (e) Fitted values of $\Gamma$, where the experimental data are fitted to the theoretical predictions with $\bar{n}=\bar{n}_{0}$. Solid line depicts $y=x$. In (d) and (e), the blue (brown) shaded region represents the expected uncertainty of the respective value when the populations are averaged over 20 (200) random trials. (f) Lorentzian spectral densities for various target (solid) and fitted (dashed) values of $\Gamma$. The shaded region represents the expected uncertainty of $\Gamma$ when averaged over 20 random trials.
First, we demonstrate the tunability of the bath’s temperature, quantified by the average phonon number $\bar{n}$ of the bath oscillator’s initial state. Resonant SDK operations with stochastically varying phase are applied to the ion in the motional ground state, which prepares the thermal state when averaged over many random trials (see Methods for details). Figure 1a illustrates the simulated coherent time evolution with respect to $\hat{H}_{SO}$, for various values of $\bar{n}$. The dynamics exhibit coherent oscillations with a slowly fluctuating envelope, where the envelope signifies resonant ($\nu=\Delta$) energy exchange between the system and the bath oscillator. This is related to the Jaynes-Cummings collapse and revival with coherent initial states Eberly et al. (1980); our initial thermal state is an ensemble of randomly displaced coherent states. A thermal state of larger $\bar{n}$ is a mixture of more “scattered” coherent states, leading to diminished revival amplitude at $t\approx 2\pi/\Delta\times 10.5$. We see a notable agreement between theoretical predictions (solid lines) and experimental results (dots).
We compare our experimental results with the predictions of noise-aware models, which our noisy experiments are expected to be simulating. We use spin-boson models with initial phonon number $\bar{n}+\bar{n}_{0}$ and Lorentzian spectral density with full width at half maximum (FWHM) $\Gamma$ [see Eq. (3)], where the values $\bar{n}_{0}=0.036$ and $\Gamma=0.0022\Delta$ are extracted from separate experiments for characterizing imperfect cooling and motional coherence time, respectively. The experimentally measured revival amplitude (Fig. 1c) and fitted $\bar{n}$ (Fig. 1d) match well with the predictions of noise-aware models. This shows that well-characterized experimental noise can be used in determining our control parameters for simulating a target model, and thus is not always a bottleneck for the accuracy of our simulations.
Next, we show the tunability of the dephasing strength, which corresponds to the linewidth of the spin-boson model’s spectral density. Dephasing is implemented by adding random frequency offsets to the SDK operations that simulate the time evolution (see Methods for detials). This mimics the random fluctuations of the motional-mode frequency, following the stochastic model of motional dephasing noise in trapped-ion systems Wang et al. (2020).
Figure 1b illustrates the dynamics for various values of $\Gamma$ between 0 and $0.4\Delta$, which exhibit a phase transition from underdamped to critically damped ($\Gamma\approx 0.18\Delta$) to overdamped dynamics as $\Gamma$ is increased. In the underdamped regime, the dephasing strength is sufficiently low such that the amplitude of oscillation revivals is diminished but not entirely vanished. We observe a phase flip occurring at the node of the envelope at $t\approx 2\pi/\Delta\times 5\>(6)$ for $\Gamma=0\>(0.05\Delta)$. Conversely, in the overdamped scenario, strong dephasing diminishes the resonant coupling between the system and the oscillator, leading to oscillations that decay indefinitely without any revivals or phase flips. Thus, after the underdamped oscillations’ revival, the phase of oscillation is opposite between the underdamped and overdamped dynamics, as clearly seen at $t\approx 2\pi/\Delta\times 10$, for example. The experimental data match well with the theoretical predictions of spin-boson models.
The uncertainty of simulated $\bar{n}$ and $\Gamma$ can be reduced by performing a larger number of random trials performed, where each trial uses a fixed set of randomly drawn parameters. In Fig. 1d and e, we perform numerical simulations comprising 2000 random trials and calculate the standard deviation of fitted values for $\bar{n}$ and $\Gamma$, which are obtained by averaging the results over either 20 or 200 trials. In our experiments, due to the control software’s delay in compiling a given set of control parameters, we choose to use 20 random trials, each trial repeated 100 times to obtain the qubit population. The experimental data indeed sit within the standard deviation for 20 trials. In future experiments with faster compilation Dalvi et al. (2023), we expect to perform a larger number of random trials with fewer repetitions each sup .
With the capability of simulating a bath with continuous spectral density, we use multiple oscillators to simulate molecular energy transfer in a model structured bath. We simulate a bath of spectral density composed of up to 3 Lorentzian peaks, where each peak represents the vibrational mode of a molecular system such as light-harvesting protein complexes Olbrich et al. (2011); Chandrasekaran et al. (2015); Lee and Coker (2016); Kim et al. (2018); Tamascelli et al. (2019). We use the electronic coupling strength $\Delta=500$ cm-1 Chenu and Scholes (2015), such that at temperature 77 K, $\bar{n}\approx 0.1$ for the near-resonance vibrational modes.
Figure 2: Simulations of molecular energy transfer in model structured baths. (a) Dashed lines represent the radial motional modes strongly coupled by the Raman beam. Black (grey) lines denote the modes with non-zero (zero) coupling strength to the target ion. The three solid lines (shaded regions) depict the mean value (standard deviation) of the randomized pulse frequencies. (b) Simulated spectral density of the spin-boson model’s bath, where the electronic coupling strength $\Delta=500\>\text{cm}^{-1}$. The dashed curves depict the Lorentzian peak generated by the pulses of corresponding colors in (a). The orange (brown) solid curves illustrate the spectral density summed over two (three) peaks. (c) Time evolution of the spin-boson models with $J(\omega)$ in (b). The bath temperature is 77 K such that $\bar{n}\approx 0.1$, which is set for trapped ions’ motional modes using imperfect sideband cooling. Curves (dots) represent theoretical predictions (experimental data). Error bars denote the standard deviation over 20 trials, each trial performed with randomly drawn set of parameters and repeated 100 times.
Figure 3: Simulations based on Leggett’s spin-boson model. (a) Spectral densities of the Leggett’s model. Dot-dashed curves are the sub-Ohmic ($s=0.5$), Ohmic ($s=1.0$), and super-Ohmic ($s=2.0$) spectral densities described by Eq. (4) ($A=0.1$, $\omega_{c}=10\Delta$). Solid curves are the spectral densities simulated in our experiments using two motional modes. These spectral densities are designed to match the dot-dashed curves near $\omega=\Delta$. Insets compare the solid and dot-dashed curves within the interval $\omega\in[0.85\Delta,1.15\Delta]$. (b) (c) (d) Dynamics of the spin-boson model for $s=0.5$, $1.0$, and $2.0$, respectively. The solid curves represent the theoretical predictions of the spin-boson models with spectral densities given by the solid curves in (a). The red dashed curves are expected results, derived from numerical simulations of the density matrix of the qubit and motional modes subject to control operations and hardware noise. Circles denote experimental data and error bars indicate the standard deviation over 60, 20, and 20 trials for (b), (c), and (d), respectively. Each trial is performed with randomly drawn set of parameters and repeated 100 times. (e) Time evolution of the spin-boson models. The solid (dot-dashed) curves show the evolution with spectral densities given by the solid (dot-dashed) curves in (a).
The spin-boson model with $J(\omega)$ in Fig. 2b is simulated using the SDK operations of pulse frequencies shown in Fig. 2a, where the pulse frequency is defined as the motion detuning of laser beams from the carrier transition. The 1st, 3rd, and 5th motional modes (in the order of increasing mode frequency) are used. Figure 2c shows the results for simulating spin-boson models with spectral densities composed of 2 and 3 Lorentzian peaks, respectively. The theoretical predictions are obtained using the time-dependent density-matrix renormalization group algorithm White and Feiguin (2004); Prior et al. (2010) in the interaction picture Nuomin et al. (2022).
The oscillations in Fig. 2c display repeated suppression and revival in amplitude, where the revival amplitude decreases with time, which is a signature of underdamped oscillations. The second suppression occurs earlier in the 3-mode curve than in the 2-mode curve, as the added Lorentzian peak increases the overall strength of spin-bath coupling. The suppressed oscillations are faster than before or after the suppression, as can be seen from the opposite oscillation phase between the 2-mode and 3-mode curves near $t\approx 650$ fs. Interestingly, the 3-mode experimental data exhibit slower oscillations during the second suppression. Nonetheless, with this exception, the experimental data overall closely align with the theoretical predictions, capturing the fine details of the oscillation envelope.
The experimental time for simulating the time evolution is 1.63 ms (2.56 ms) for the 2-mode (3-mode) experiment, which is in the same order of magnitude as the motional coherence time $5.18(67)$ ms that characterizes motional dephasing, the leading source of errors in our trapped-ion system Kang et al. (2023a). Note that motional dephasing, when modelled as stochastic fluctuations of motional-mode frequencies Wang et al. (2020), simply increases the simulated value of $\Gamma$. A simulation of coherent spin-oscillator model with the same experimental time would have substantial errors due to motional dephasing Kang et al. (2024). However, for simulations of dephased spin-oscillator models, motional dephasing may have negligible effects, as the increase in $\Gamma$ may be small relative to the target value. To improve the accuracy of simulations with longer experimental time, the standard deviation of random frequency offsets may be reduced such that the effects of motional dephasing are compensated. This extends the capability of noisy hardware for simulating open quantum systems by essentially using the noise Guimarães et al. (2023); Leppäkangas et al. (2023). Reliable utilization of noise may require careful characterization, such as spectroscopy of motional dephasing Kang et al. (2023a); Milne et al. (2021); Keller et al. (2021).
We further simulate the spin-boson models inspired by Leggett et al. Leggett et al. (1987). The spectral density is given by
$J_{\rm Legg}(\omega)=A\omega^{s}\omega_{c}^{1-s}e^{-\omega/\omega_{c}},$ (4)
where $s<1$, $s=1$, and $s>1$ represent the sub-Ohmic, Ohmic and super-Ohmic baths, respectively, and $\omega_{c}$ is the cutoff frequency. This model is widely used to characterize noise in nanomechanical devices Seoanez et al. (2007); Galland et al. (2008); Chen et al. (2011), superconducting circuits Magazzù et al. (2018); Thorwart et al. (2004); Kaur et al. (2021); Gulácsi and Burkard (2023), and proton-transfer reactions Cukier and Morillo (1989); Ohta et al. (2006); Craig et al. (2007). The spin dynamics are expected to exhibit a phase transition from coherent oscillations to incoherent decay or localization as $s$ is decreased or $A$ is increased, which is a subject of longstanding research using analytical tools Leggett et al. (1987); Bulla et al. (2003); Vojta et al. (2005) and classical simulations Wang and Thoss (2008); Duan et al. (2017); Strathearn et al. (2018); Kast and Ankerhold (2013); Nalbach and Thorwart (2013); Otterpohl et al. (2022); Nacke et al. (2023). Adding quantum simulators to the list of methods for studying this model is an interesting task.
We use a sum of Lorentzian spectral densities Meier and Tannor (1999) to approximate the spectral density $J_{\rm Legg}(\omega)$. To find the optimal set of parameters $\kappa$’s, $\Gamma$’s, and $\nu$’s in Eq. (3) that best represent $J_{\rm Legg}(\omega)$, well-established global optimization algorithms Wolz (2022) are used to minimize the cost function, which is defined as the integral of the absolute difference between $J_{\rm Legg}(\omega)$ and the sum of corresponding Lorentzian spectral densities over a given bandwidth sup .
As a first step toward simulating Leggett’s spin-boson models, we consider weak spin-bath coupling ($A=0.1$) and minimize the cost function for a limited bandwidth near resonance $\omega\in[0.9\Delta,1.1\Delta]$ using 2 Lorentzian peaks. Figure 3a shows the obtained sums of Lorentzian peaks (solid lines) that approximate sub-Ohmic ($s=0.5$), Ohmic ($s=1.0$), and super-Ohmic ($s=2.0$) spectral densities (dot-dashed lines), respectively. The results for simulating the spin-boson models with these approximated spectral densities are shown in Fig. 3b-d. We observe a crossover from near-coherent ($s=2$) to fully damped ($s=0.5$) oscillations in the simulated time scale. The experimental data match well with theoretical predictions, where the deviations in the $s=2$ case result from insufficient number of Trotterization steps sup .
Figure 4: Simulating sub-Ohmic ($s=0.5$) bath with 10 oscillators. (a) Comparison between the sub-Ohmic spectral density (dot-dashed blue) and its approximation constructed by 10 noisy oscillators (solid red). Each dashed line represents the Lorentzian spectral density generated by each of these 10 oscillators. (b) Comparison of the population dynamics generated by the sub-Ohmic spectral density (dot-dashed blue) and its 10-mode approximation (solid red).
Figure 3e compares the predicted population dynamics of the spin-boson model with spectral density $J_{\rm Legg}(\omega)$ and the corresponding approximated spectral density. While the dynamics match very well for the Ohmic and super-Ohmic cases, the sub-Ohmic case shows a non-negligible deviation. This is because the $J(\omega)$ values are overall larger for the sub-Ohmic bath, so matching the population dynamics requires accurately capturing a larger bandwidth. Indeed, in future experiments with a larger number of motional modes, we expect to match a wider range of spectral density and thus more accurately simulate the sub-Ohmic spin-boson model, as suggested by the 10-mode example in Fig. 4. Note that Ref. Porras et al. (2008) suggests using roughly 100 motional modes to represent the Ohmic spectral density. Using noisy modes allows simulating the spin-boson models much more efficiently in terms of the number of motional modes used.
Finally, we simulate the VAET model Gorman et al. (2018) with nonzero energy detuning $\epsilon$ between the spin states. In the absence of coupling to a vibrational mode ($\kappa=0\>\text{cm}^{-1}$), energy transfer between the detuned spin states is suppressed; however, for nonzero $\kappa$, the energy transfer can be activated if the mode frequency $\nu$ meets the resonance condition. Here we set the spin parameters as $\Delta=30\>\text{cm}^{-1}$ and $\epsilon=100\>\text{cm}^{-1}$.
Figure 5: Simulations of vibrationally-assisted energy transfer (VAET). $\epsilon=100\>\text{cm}^{-1}$ and $\Delta=30\>\text{cm}^{-1}$ are fixed. (a) (b) Time evolution of the donor-state population for coherent VAET models with $\kappa=30\>\text{cm}^{-1}$. Solid (dashed) curves depict theoretical predictions of the target (noise-aware) model. Circles represent experimental data and error bars are standard deviation over 100 repetitions. In (a), bath temperature is fixed to 0 K ($\bar{n}=0$) and $\nu$ is varied. In (b), $\nu$ is fixed to the resonant value $104\>\text{cm}^{-1}$ and two temperatures 0 K and 300 K, which correspond to $\bar{n}=0$ and 2.0, respectively, are simulated. (c) Time evolution of the donor-state population for noisy VAET models with $\nu=104\>\text{cm}^{-1}$ and $\Gamma=10\>\text{cm}^{-1}$ at 0 K. Solid (dashed) curves depict theoretical predictions of the spin-boson (dephased spin-oscillator) model. Circles denote experimental data. For $\kappa=0\>\text{cm}^{-1}$, error bars indicate the standard deviation over 100 repetitions. For $\kappa=10$ and $20\>\text{cm}^{-1}$, error bars indicate the standard deviation over 20 trials, each trial performed with randomly drawn set of parameters and repeated 100 times.
We first explore the coherent spin-oscillator model $\hat{H}_{SO}$ with a relatively large coupling strength $\kappa=30\>\text{cm}^{-1}$ and various values of $\nu$. When $\nu$ satisfies the resonance condition of VAET, i.e., $\nu=\sqrt{\Delta^{2}+\epsilon^{2}}/k\approx 104\>\text{cm}^{-1}/k$ where $k$ is an integer, energy transfer occurs Gorman et al. (2018). Otherwise, energy remains localized at the donor site. Figure 5a shows the experimental data for simulating this model for the resonant ($\nu=52$ and $104\>\text{cm}^{-1}$) and non-resonant ($\nu=78\>\text{cm}^{-1}$) cases.
Furthermore, in Fig. 5b, we simulate the impact of initial temperature on the resonant ($\nu=104\>\text{cm}^{-1}$) VAET. We simulate baths of two temperatures 0 K and 300 K, which correspond to initial average phonon number $\bar{n}=0$ and $2.0$, respectively. For the latter case, SDK operations with random phase are applied to prepare the thermal state (when averaged over many trials). At first, the energy transfer is faster for the case of higher temperature, as more vibrational quanta assist the energy transfer. However, at later times, high temperature leads to a rapid decay of oscillations, as a thermal state of large $\bar{n}$ is a highly mixed state which quickly loses the coherent behavior. This pattern is accurately captured by our trapped-ion simulations.
Noise (dephasing or damping) can be added to the VAET model. In Fig. 5c we simulate the dephased spin-oscillator model with $\Gamma=10\>\text{cm}^{-1}$ and various values of $\kappa$, showing good match with theoretical predictions. However, as $\kappa$ is increased, the results deviate more from the theoretical predictions of the damped spin-oscillator model, which is non-perturbatively equivalent to those of the spin-boson model with $J(\omega)=J_{\rm Lo}(\omega)$ Lemmer et al. (2018); Tamascelli et al. (2018). This aligns with the aforementioned expectation that our method of using the dephased spin-oscillator model to approximate the spin-boson model is valid only in the weak-$\kappa$ regime. To summarize, the short-time dynamics between the two models match as the 2-point correlation functions are the same; however, the long-time dynamics deviate due to mismatch in the higher-order correlation functions sup . Indeed, for nonzero $\epsilon$, the steady-state donor populations are different between the dephased spin-oscillator model and the spin-boson model, as clearly shown by the $\kappa=20\>\text{cm}^{-1}$ curves.
Discussion and outlook
An immediate future direction is to observe the effects of noisy modes on the signature behaviors of small-scale molecular models using trapped-ion simulations Kang et al. (2024). Examples are: conical intersections Whitlow et al. (2023); Valahu et al. (2023), polarized-light-induced electron transfer Sun et al. (2023); Zhang et al. (2024), and VAET in the donor-bridge-acceptor system Li et al. (2021). This may involve performing operations on a qudit Sun et al. (2023) or more than one qubits. For the latter case, entangling operations between the qubits Mølmer and Sørensen (1999) need to be added to the Trotterized instruction of operations.
Our method of using randomized operations to simulate spin-boson models is limited to weak-coupling regime. However, simulating a bath of dephased oscillators may be of independent interest, as this can be thought as a natural extension from Gaussian ($\hat{L}\propto\hat{b}$ or $\hat{b}^{\dagger}$) to non-Gaussian ($\hat{L}\propto\hat{b}^{\dagger}\hat{b}$) bath models Linowski et al. (2022); Barthel and Zhang (2022). In such case, strong-coupling regime may be readily explored. For example, an interesting question is whether the phase transition from coherent oscillations to localization for sub-Ohmic spin-boson models Nalbach and Thorwart (2013); Duan et al. (2017); Otterpohl et al. (2022) can be observed using dephased spin-oscillator models that match the spectral densities. As evidenced in Fig. 4, this may be simulated by our method applied to a few tens of motional modes, which can be readily utilized in a chain of 10 to 20 ions.
Simulations of spin-boson models in strong-coupling regime can be performed using the damped spin-oscillator model, as its spin dynamics are non-perturbatively equivalent to those of the spin-boson model with the same spectral density Lemmer et al. (2018); Tamascelli et al. (2018). Damped oscillators can be realized using sympathetic cooling of the trapped ions’ motional modes Lemmer et al. (2018); Schlawin et al. (2021). This requires using a multi-species ion chain, such that the optical pumping during sympathetic cooling does not destroy the qubit states that encode the spins. If accomplished, this implementation of damping can be combined with our dephasing method to simulate unexplored models of open quantum systems. We note that while the role of electronic states’ dephasing in molecular energy transfer has been heavily studied Grad et al. (1988); Goldsmith et al. (2006); Plenio and Huelga (2008); Hayes et al. (2010); Moreira et al. (2020), the role of vibrational modes’ dephasing is relatively less explored Yoneda et al. (2021).
An ultimate goal is to accurately simulate the quantum dynamics of various models in physical Caldeira and Leggett (1983); Weiss et al. (1987), chemical Garg et al. (1985); Gilmore and McKenzie (2005), and biological Huelga and Plenio (2013); Jang and Mennucci (2018) systems that are intractable using classical simulation methods MacDonell et al. (2021); Kang et al. (2024). Classical simulation approaches continue to improve with advancements in numerical methods, yet they encounter fundamental challenges when dealing with large quantum systems. These challenges stem from the exponentially large Hilbert space and the quantum entanglement between the components. Take, for example, the Dissipation-Assisted Matrix Product Factorization (DAMPF) method Somoza et al. (2019), a state-of-the-art method for simulating open quantum systems. For simulating a one-dimensional chain of $N_{s}$ electronic states, each coupled to a bath of $N_{b}$ damped bosonic modes, the computational complexity of DAMPF scales as $\mathcal{O}(N_{s}^{5}N_{b}D_{b}^{2}\chi^{3})$, where $D_{b}$ is the number of energy levels that represent each bosonic mode and $\chi$ is the bond dimension in the matrix product operator representation of the system-bath density matrix Somoza et al. (2019). A critical issue is the rapid, potentially exponential increase in $\chi$ with respect to the system-bath coupling strength, for a fixed target accuracy in the population dynamics. In comparison, trapped-ion simulators may offer a more scalable alternative, as the number of operations scales as only $\mathcal{O}(N_{s}N_{b})$ and the total duration of SDK operations grows linearly with the system-bath coupling strength Kang et al. (2024). Also note that classical simulation methods typically struggle with complicated connectivity between the components, while for trapped ions all-to-all connectivity is granted both between qubits and between qubits and motional modes.
Realizing accurate trapped-ion simulations of models with many spins and noisy bosonic modes requires overcoming more challenges, such as efficiently characterizing the motional-mode parameters Kang et al. (2023b) and significantly lowering the hardware’s noise level. Nonetheless, our demonstration of accurate simulations using controlled dissipation represents a key step forward in addressing these challenges. In particular, the accuracy of our results is enabled by carefully designed calibration schemes and precise phase tracking sup ; Lobser et al. (2023). Also, our method enables using motional dephasing noise for simulating open quantum systems, which significantly relaxes the limitation on the experiment time (for simulating the time evolution) of our simulations imposed by hardware noise.
We anticipate this work inspires future works on improving the scale and accuracy of programmable quantum simulations of open quantum systems. Furthermore, our approach of finding models and parameter regimes that allow utilizing the dominant hardware noise may be of interest for various other analog quantum simulation platforms.
During the preparation of our manuscript we became aware of Ref. So et al. (2024) which simulates the electron-transfer model using sympathetic cooling on the trapped ions’ motional mode. This complements our experiment on simulating spin-boson models with structured bath using randomized operations.
Acknowledgements.
The authors thank Yichao Yu for his valuable discussions about the experimental setup, Daniel Lobser for his valuable suggestion in controlling the RFSoC control software, and Thomas Barthel for his course on open quantum systems that inspired this work. This research is funded by the Office of the Director of National Intelligence - Intelligence Advanced Research Projects Activity, through the ARO contract W911NF-16-1-0082 (provided the experimental apparatus used in the demonstration), the DOE BES award DE-SC0019400 (theoretical modeling), and the NSF Quantum Leap Challenge Institute for Robust Quantum Simulation Grant No. OMA-2120757 through a seed grant (experimental concepts and validation). Support is also acknowledged from the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator (data analysis and validation).
Methods
Experimental implementation
Figure 6: Experimental setup and instruction of operations. (a) Cartoon depiction of the experimental setup illustrates the orientation of components. The magnetic field is perpendicular to the trap surface, while the individual beam is perpendicular to the trap surface and the global beam is parallel to it. The differential wave vector $\Delta\textbf{k}$ is oriented at a 45-degree angle to the trap surface, aligning with one of the radial axes. The global beam comprises two tones driving the blue (BSB) and red (RSB) sideband transitions, respectively. (b) Simplified energy levels of the 171Yb+ ion and the laser pulses used for implementing the on-resonance spin-dependent kick (SDK) operation. (c) Visualization of the motional-state evolution in phase space when a SDK operation of duration $\tau$ is applied, where the qubit is in the eigenstate of $\hat{\sigma}_{\phi_{s}}\equiv\hat{\sigma}_{+}e^{i\phi_{s}}+\hat{\sigma}_{-}e^{-i\phi_{s}}$. Here, $\hat{\sigma}_{+}$ ($\hat{\sigma}_{-}$) is the raising (lowering) operator of the qubit, $\phi_{s}$ ($\phi_{m}$) is the spin (motion) phase of the laser beams, $\tilde{\Omega}$ is the Rabi frequency of the sideband transition, and $\delta_{m}$ is the motion detuning. (d) Instruction of operations for simulating the dephased spin-oscillator model, which approximates the spin-boson model. ‘H’ represents the Hadamard gate on the qubit. ‘Kick (heat)’ represents the SDK operations with random $\phi_{m}$ that prepare the thermal state on average. ‘SQ’ (‘Kick (coh/dep)’) represents the single-qubit rotations (SDK operations) that simulate the first (second) term in Eq. (5). Coherent (dephased) spin-oscillator model is simulated by using a fixed (randomly drawn) $\delta_{m}$ value at each step.
The simulations of spin-boson models are performed using a linear chain of 171Yb+ ions confined in a micro-fabricated surface trap Revelle (2020), as illustrated in Fig. 6a. The spin states are encoded to the 2S1/2-orbital hyperfine energy levels (qubit states) as |$\ket{0}\equiv\ket{F=1;m_{F}=0}$ and |$\ket{1}\equiv\ket{F=0;m_{F}=0}$. The bosonic bath modes are encoded as the collective radial motional modes of the ion chain. Details of our experimental setup can be found in Ref. Wang et al. (2020).
Through the sequential application of Doppler cooling, EIT cooling, and sideband cooling techniques, the zigzag motional mode of the ion chain can be cooled to $\bar{n}=0.036(16)$. Following this, a $370$ nm laser is utilized to pump the ions to their ground state |$\ket{1}$. Transitions between the qubit states are performed using stimulated Raman transitions, driven by a pair of laser beams Wang et al. (2020). Acousto-optic modulators are employed to manipulate the frequency and phase of each laser beam, thereby enabling the execution of Raman transitions at specified frequency detunings and phases. Specifically, the spin-oscillator coupling is simulated using the spin-dependent kick (SDK) operation induced by simultaneous blue- and red-sideband transitions as shown in Fig. 6b. Figure 6c visualizes how the motional state evolves in phase space when a SDK operation of duration $\tau$, spin (motion) phase $\phi_{s}$ ($\phi_{m}$), motion detuning $\delta_{m}$ (see Fig. 6b), and sideband Rabi frequency $\tilde{\Omega}$ is applied Lee et al. (2005). Subsequent to the final operation, the qubit-state population is accessed via the state-dependent detection technique.
We strategically use the center ion of a 7-ion chain, as it exhibits negligible coupling to the 2nd, 4th, and 6th modes (See Fig. 7). This arrangement maximizes the effective spacing between the motional-mode frequencies and therefore suppresses cross-mode coupling (driving modes that are not targeted by the SDK operation).
Figure 7: Ion-Mode coupling strengths of a 7-ion chain. Motional-mode frequencies are shown in the lower right corner of each panel.
Trotterization
The evolution of $\hat{H}_{SO}$ in Eq. (2) is mapped to a sequence of single-qubit rotations and SDK operations via Trotterization. After conjugating $\hat{H}_{SO}$ with a Hadamard gate on the qubit, the Hamiltonian can be described in the interaction picture as
$\hat{H}_{I}(t)=\frac{\epsilon}{2}\hat{\sigma}_{\Delta}(t)+\frac{\kappa}{2}\hat{\sigma}_{\Delta}(t)(\hat{b}e^{-i\nu t}+\hat{b}^{\dagger}e^{i\nu t}),$ (5)
where $\hat{\sigma}_{\Delta}(t)\equiv\hat{\sigma}_{+}e^{i\Delta t}+\hat{\sigma}_{-}e^{-i\Delta t}$ and $\hat{\sigma}_{+}$ ($\hat{\sigma}_{-}$) is the raising (lowering) operator of the qubit. The evolution with respect to $\hat{H}_{I}(t)$ is discretized into $N$ time steps, each consisting of a single-qubit rotation and a SDK operation that simulate the first and second term, respectively. Specifically, for the SDK operation at the $j$-th time step, the spin (motion) phase of the laser beams Lee et al. (2005) is set as $\Delta t_{j}$ ($\nu t_{j}$), where $t_{j}\equiv(j-\frac{1}{2})T/N\tau$ and $\tau$ is the duration of the operation for each time step. This is equivalent to setting the symmetric (anti-symmetric) detuning from the sideband transitions, referred to as the spin (motion) detuning, as $\Delta T/N\tau$ ($\nu T/N\tau$) at all time steps. This method can be straightforwardly extended to the case where multiple modes are used; however, the laser phase needs to be tuned to track the phase of each motional mode correctly sup .
The Trotterization method here enables simulating the VAET model on a single qubit. This is different from Ref. Gorman et al. (2018), which simulates the same model using continuous evolution of two ions induced by global laser beams. Thus, Trotterization using individually addressed beams is more efficient in terms of the required number of qubits. Also, our approach can be straightforwardly extended to simulating more complicated molecular models consisting of many electronic states and vibrational modes Kang et al. (2024).
Randomized operations
Randomizing the control parameters is a key ingredient of our simulations of spin-boson models. By taking an average of many coherent evolutions, each involving a stochastically varying parameter, we realize both heating and dephasing as described by the Lindblad master equation Cai and Barthel (2013); Chenu et al. (2017). Heating (dephasing) is used to tune the temperature (spectral linewidth) of the bath.
First, infinite-temperature heating, described by the pair of Lindblad operators $\hat{L}_{1}=\sqrt{\Gamma^{\prime}}\hat{b}$ and $\hat{L}_{2}=\sqrt{\Gamma^{\prime}}\hat{b}^{\dagger}$, is induced by applying SDK operations with stochastically varying motion phase $\phi_{m}$. Specifically, we apply $N^{\prime}$ resonant ($\delta_{m}=0$) SDK operations of duration $\tau^{\prime}=4\Gamma^{\prime}/\tilde{\Omega}^{2}$, spin phase $\phi_{s}=0$, and $\phi_{m}$ randomly drawn from $[0,2\pi)$ for each operation. These stochastic operations are applied on the initial qubit-mode composite state |$\ket{+,n=0}$; note that $\ket{+}$ is an eigenstate of the spin operator with $\phi_{s}=0$, and thus the qubit is decoupled from these operations. When averaged over many trials of coherent evolutions, each executed with a distinct set of $N^{\prime}$ random $\phi_{m}$ values, this procedure prepares the thermal state with an average phonon number $\bar{n}=N^{\prime}\Gamma^{\prime}\tau^{\prime}$. This can be intuitively understood as the thermal state represented as an ensemble of randomly displaced coherent states. See the Supplemental Material sup for derivations.
For the experimental data in Fig. 1a and Fig. 5b, the average phonon number $\bar{n}$ of the initial thermal state is tuned by varying the duration $\tau^{\prime}$ of SDK operations, where the number of steps $N^{\prime}$ is fixed to 5. The errors in $\bar{n}$ due to finite number of steps is discussed in the Supplemental Material sup .
Next, dephasing during the time evolution, which is described by a single Lindblad operator $\hat{L}_{1}=\sqrt{\Gamma}\hat{b}$, is induced by giving random offsets to the motion detuning $\delta_{m}$ of the $N$ Trotterized SDK operations. Specifically, an offset to $\delta_{m}$ is assigned for each of the $N$ steps, with values drawn from a normal distribution characterized by a mean of zero and a standard deviation of $\sqrt{\Gamma T/N}/\tau$ sup . The qubit population dynamics averaged over many trials of coherent evolutions, each performed with a set of $N$ random $\delta_{m}$ values, give the dynamics with respect to the dephased spin-oscillator model. The entire instruction of operations is shown in Fig. 6d.
In practice, the dephasing strength $\Gamma$ is varied by tuning the standard deviation of the random offsets to $\delta_{m}$ at each Trotterization step. This enables tuning the FWHM of each Lorentzian peak in the bath’s spectral density. The errors in the simulated dephasing strength due to finite number of Trotterization steps is discussed in the Supplemental Material sup .

### v2

@@
 Abstract
-The spin-boson model, involving a spin interacting with a bath of quantum harmonic oscillators, is a widely used representation of open quantum systems.
-Trapped ions present a natural platform for the quantum simulation of such models; however, previous experiments have been limited to simulating a scenario where a spin is coherently coupled to bosonic modes, which neglects capturing the dissipation of the bath.
-In our work, we perform quantum simulations of spin-boson models with structured baths using the motional states of trapped ions.
-We demonstrate the capability for adjusting the bath’s temperature and continuous spectral density by adding randomness to the control parameters.
-Subsequently, we simulate the dynamics of various spin-boson models with spectral densities composed of up to three Lorentzian peaks.
-The experimental outcomes closely align with theoretical predictions, suggesting the advantage of adding the trapped-ion system’s experimental noise for simulating open quantum systems.
-Every quantum system inevitably interacts with the surrounding environment.
-Such interactions can significantly influence the dynamics of the system.
+The spin-boson model, involving spins interacting with a bath of quantum harmonic oscillators, is a widely used representation of open quantum systems.
+Trapped ions present a natural platform for simulating the quantum dynamics of such models, thanks to the presence of both high quality internal qubit states and the motional modes of the ions that can simulate the relevant quantum degrees of freedom.
+In our work, we extend the previous body of work that focused on coherent coupling of the spins and bosons to perform quantum simulations with structured dissipative baths using the motional states of trapped ions.
+We demonstrate the capability for adjusting the bath’s temperature and continuous spectral density by adding randomness to fully programmable control parameters.
+Subsequently, we simulate the dynamics of various spin-boson models with noise spectral densities constructed from coupling to several dissipative harmonic oscillator modes.
+The experimental outcomes closely align with theoretical predictions, indicating successful simulation of open quantum systems using a trapped-ion system.
+Interaction of a quantum system with its environment in an open dissipative system can significantly influence the dynamics of the system.
 A paradigmatic model of non-Markovian open quantum systems that captures the quantum nature of both the system and the environment is the spin-boson model Leggett et al.
 (1987), which is ubiquitous in the study of condensed-matter physics Caldeira and Leggett (1983); Weiss et al.
+(1987); Schindler et al.
 (2022), chemical reactions Garg et al.
-(1985); Gilmore and McKenzie (2005), and biological light-harvesting complexes Huelga and Plenio (2013); Jang and Mennucci (2018).
+(1985); Gilmore and McKenzie (2005), collective modes in atomic-photonic systems Strack and Sachdev (2011); Bhaseen et al.
+(2012), and biological light-harvesting complexes Huelga and Plenio (2013); Jang and Mennucci (2018).
 Experimentally realizing the dynamics of spin-boson models in a programmable fashion is an interesting challenge for quantum simulators Mostame et al.
@@
 The bosonic bath can be simulated using either a large number ($\gtrsim 50$) of motional modes Porras et al.
-(2008) or a few dissipative motional modes Lemmer et al.
+(2008) or a handful of dissipative motional modes Lemmer et al.
 (2018); Schlawin et al.
@@
 (2022); Chen et al.
-(2023), past works in this direction are limited to simulating a spin either coherently coupled to up to two bosonic modes Gorman et al.
+(2023a), past works in this direction are limited to simulating a spin coherently coupled to up to two bosonic modes Gorman et al.
 (2018); MacDonell et al.
@@
 (2023).
-These previous studies did not consider the effects of dissipation, which is often described by the coupling of a system to a continuum of bath modes.
-Dissipation of the environment determines the system’s long-term behavior, such as reaching a steady state, and therefore is essential to incorporate for accurately modeling open quantum systems.
-Note that a recent study Wang et al.
-(2024) simulated a spin dissipatively coupled to many modes albeit with limited tunability.
+These previous studies did not consider the effects of dissipation, which is often described by the coupling of a system to a continuum of bath modes, that can determine the system’s long-term behavior such as reaching a steady state.
 In this work, we perform quantum simulations of spin-boson models with continuous spectral densities using trapped ions and their motional modes.
@@
 The various spin-state population dynamics, which display behaviors such as damped oscillations, coherent revivals, and relaxation to a steady state, are accurately captured by our quantum-simulation results.
-Capturing the dissipative behavior of the spin-boson model, as well as tuning the bath’s temperature, is enabled by adding randomness to the control parameters such as laser frequency and phase.
-Remarkably, adding random frequency offset to the control of trapped ions’ motional modes is directly analogous to the leading source of noise of our experimental setup, namely, motional dephasing Kang et al.
+We note that a recent study Wang et al.
+(2024) simulated a spin dissipatively coupled to many modes, albeit with limited tunability.
+Capturing the dissipative behavior of the spin-boson model, as well as tuning the bath’s temperature, is enabled by adding randomness to the control parameters, such as the laser frequency and phase.
+Remarkably, the impact of these control signals is directly analogous to motional dephasing, the leading source of noise in our experimental setup Kang et al.
 (2023a).
-Thus, motional dephasing noise of the hardware can essentially be used by our quantum simulations, providing accurate results even when the experimental time for simulating the time evolution is comparable to the coherence time of the motional modes.
+This serves as a lower limit for the level of noise that can be introduced in our quantum simulation process.
 While the quantum simulations presented here are limited to the weak spin-bath coupling regime, our demonstration of adding or using noise Guimarães et al.
 (2023); Leppäkangas et al.
-(2023) in analog quantum simulations is an important ingredient for potential quantum advantage in simulating open quantum system models that are intractable using classical simulation methods MacDonell et al.
+(2023) in analog quantum simulations is an important ingredient for a potential quantum advantage in simulating open quantum system models that are intractable using classical simulation methods MacDonell et al.
 (2021); Kang et al.
@@
 Models
-The spin-boson model is described by a spin coupled to a continuous bath of quantum harmonic oscillators. The Hamiltonian is given by ($\hbar$ is set as $1$ for simplicity)
+The spin-boson model describes a spin coupled to a continuous bath of quantum harmonic oscillators. The Hamiltonian is given by
 […]
 […]
-A spin-boson model with structured bath is often characterized by a spectral density $J(\omega)$ composed of multiple Lorentzian peaks, each of which can be represented by a noisy harmonic oscillator Lemmer et al.
+A spin-boson model with a structured bath is often characterized by a spectral, density $J(\omega)$, composed of multiple Lorentzian peaks, each of which can be represented by a noisy harmonic oscillator Lemmer et al.
 (2018).
@@
 $J_{\rm Lo}(\omega)=\kappa^{2}\Big{(}\frac{\Gamma/2}{(\Gamma/2)^{2}+(\omega-\nu)^{2}}-\frac{\Gamma/2}{(\Gamma/2)^{2}+(\omega+\nu)^{2}}\Big{)},$ (3)
-where we assume that initially the spin is in the |$\ket{0}$ state and the oscillator is in the thermal state of average phonon number $\bar{n}(\nu)$.
-The spectral density of the bath determines the dynamics of the spin up to leading order in $\kappa T$, where $T$ is the evolution time Breuer and Petruccione (2002); sup .
-Therefore, the spin dynamics of the dephased spin-oscillator model approximately match those of the spin-boson model with $J(\omega)=J_{\rm Lo}(\omega)$ in the weak-$\kappa$ regime, which straightforwardly extends to the model with multiple oscillators.
-In particular, for the unbiased ($\epsilon=0$) spin-boson model, which is of particular interest Leggett et al.
+where we assume that the oscillator is initially in the thermal state with average phonon number $\bar{n}(\nu)$.
+The spectral density of the bath determines the dynamics of the spin to leading order in $\kappa T$, where $T$ is the evolution time Breuer and Petruccione (2002); sup .
+Therefore, the spin dynamics of the dephased spin-oscillator model approximately match those of the spin-boson model with $J(\omega)=J_{\rm Lo}(\omega)$ in the weak-$\kappa$ regime, which extends straightforwardly to the model with multiple oscillators.
+Of particular interest is the unbiased ($\epsilon=0$) spin-boson model Leggett et al.
 (1987); Wang and Thoss (2008); Duan et al.
 (2017); Strathearn et al.
-(2018), the equilibrium donor-state (|$\ket{0}$) population is $1/2$ for both dephased spin-oscillator models and spin-boson models.
+(2018), where the equilibrium donor-state population is $1/2$ for both dephased spin-oscillator models and spin-boson models.
 Thus, for weak spin-bath coupling, the population dynamics of the two models show an excellent match at all times; the match in short (long) time scale is achieved due to matching spectral densities (equilibrium populations).
@@
 (2018); sup .
-However, realizing damped oscillators on trapped ions requires sympathetic cooling on a chain of multiple species of ions Lemmer et al.
+Damped oscillators can be realized in trapped ions using sympathetic cooling on a chain of ions with multiple atomic species or isotopes Lemmer et al.
 (2018); Schlawin et al.
-(2021), which is challenging to perform together with coherent operations that simulate the time evolution.
-Thus, we instead use dephased oscillators to represent the structured bath, as dephasing can be readily performed using randomized operations on trapped ions as described below.
-Experimental results
+(2021).
+In this work, we focus on dephased oscillators to represent the structured bath, which is readily performed using programmable randomized control operations on a single species of trapped ions.
+Single oscillator model
 The dephased spin-oscillator model can be simulated using trapped ions by applying spin-dependent kick (SDK) operations, which induce the coupling between the trapped-ion qubit and the motional mode (see Methods for details).
-Also, by adding randomness to the control parameters of the SDK operations, we can (i) prepare the thermal state and (ii) simulate the dephasing of the motional mode, by averaging the results over many random trials.
-The key idea is that noisy (heated or dephased) evolution, described by the Lindblad master equation, is equivalent to an average of many coherent evolutions, each with respect to a stochastic Hamiltonian Cai and Barthel (2013); Chenu et al.
+By adding randomness to the control parameters of the SDK operations, we can also prepare the thermal state and simulate the dephasing of the motional mode, by averaging the results over many random trials.
+The key idea is that noisy (heated or dephased) evolution, described by the Lindblad master equation, is equivalent to an average of many coherent evolutions, each subject to a stochastic Hamiltonian Cai and Barthel (2013); Chenu et al.
 (2017).
@@
 The shaded region represents the expected uncertainty of $\Gamma$ when averaged over 20 random trials.
-First, we demonstrate the tunability of the bath’s temperature, quantified by the average phonon number $\bar{n}$ of the bath oscillator’s initial state.
-Resonant SDK operations with stochastically varying phase are applied to the ion in the motional ground state, which prepares the thermal state when averaged over many random trials (see Methods for details).
+We first demonstrate the programmability of the oscillator bath’s temperature, quantified by the average phonon number $\bar{n}$ of the bath oscillator’s initial state.
+Controlled amounts of resonant SDK operations with stochastically varying phase are applied to the ion in the motional ground state, which prepares the thermal state when averaged over many random trials (see Methods for details).
 Figure 1a illustrates the simulated coherent time evolution with respect to $\hat{H}_{SO}$, for various values of $\bar{n}$.
-The dynamics exhibit coherent oscillations with a slowly fluctuating envelope, where the envelope signifies resonant ($\nu=\Delta$) energy exchange between the system and the bath oscillator.
-This is related to the Jaynes-Cummings collapse and revival with coherent initial states Eberly et al.
-(1980); our initial thermal state is an ensemble of randomly displaced coherent states.
-A thermal state of larger $\bar{n}$ is a mixture of more “scattered” coherent states, leading to diminished revival amplitude at $t\approx 2\pi/\Delta\times 10.5$.
+The dynamics exhibit coherent oscillations with an envelope of slow beat-note that signifies resonant ($\nu=\Delta$) energy exchange between the system and the bath oscillator Eberly et al.
+(1980).
+Our initial thermal state is prepared as an ensemble of randomly displaced coherent states.
+A thermal state of larger $\bar{n}$ is a mixture of more “scattered” coherent states, leading to reduced beat-note amplitude at $t\approx 2\pi/\Delta\times 10.5$.
 We see a notable agreement between theoretical predictions (solid lines) and experimental results (dots).
-We compare our experimental results with the predictions of noise-aware models, which our noisy experiments are expected to be simulating.
+We compare our experimental results with the predictions of noise-aware models, reflecting the noise processes in the experimental setup.
 We use spin-boson models with initial phonon number $\bar{n}+\bar{n}_{0}$ and Lorentzian spectral density with full width at half maximum (FWHM) $\Gamma$ [see Eq.
-(3)], where the values $\bar{n}_{0}=0.036$ and $\Gamma=0.0022\Delta$ are extracted from separate experiments for characterizing imperfect cooling and motional coherence time, respectively.
+(3)], where the values $\bar{n}_{0}=0.036$ and $\Gamma=0.0022\Delta$ are extracted from separate experiments to characterize imperfect cooling and finite motional coherence time, respectively.
 The experimentally measured revival amplitude (Fig.
@@
 1d) match well with the predictions of noise-aware models.
-This shows that well-characterized experimental noise can be used in determining our control parameters for simulating a target model, and thus is not always a bottleneck for the accuracy of our simulations.
+This shows that well-characterized experimental noise can be used as part of the dissipation model, and thus is not always a bottleneck for the accuracy of our simulations.
 Next, we show the tunability of the dephasing strength, which corresponds to the linewidth of the spin-boson model’s spectral density.
-Dephasing is implemented by adding random frequency offsets to the SDK operations that simulate the time evolution (see Methods for detials).
+Dephasing is implemented by adding random frequency offsets to the SDK operations that simulate the time evolution (see Methods for details).
 This mimics the random fluctuations of the motional-mode frequency, following the stochastic model of motional dephasing noise in trapped-ion systems Wang et al.
 (2020).
-Figure 1b illustrates the dynamics for various values of $\Gamma$ between 0 and $0.4\Delta$, which exhibit a phase transition from underdamped to critically damped ($\Gamma\approx 0.18\Delta$) to overdamped dynamics as $\Gamma$ is increased.
-In the underdamped regime, the dephasing strength is sufficiently low such that the amplitude of oscillation revivals is diminished but not entirely vanished.
-We observe a phase flip occurring at the node of the envelope at $t\approx 2\pi/\Delta\times 5\>(6)$ for $\Gamma=0\>(0.05\Delta)$.
-Conversely, in the overdamped scenario, strong dephasing diminishes the resonant coupling between the system and the oscillator, leading to oscillations that decay indefinitely without any revivals or phase flips.
-Thus, after the underdamped oscillations’ revival, the phase of oscillation is opposite between the underdamped and overdamped dynamics, as clearly seen at $t\approx 2\pi/\Delta\times 10$, for example.
-The experimental data match well with the theoretical predictions of spin-boson models.
-The uncertainty of simulated $\bar{n}$ and $\Gamma$ can be reduced by performing a larger number of random trials performed, where each trial uses a fixed set of randomly drawn parameters.
-In Fig.
-1d and e, we perform numerical simulations comprising 2000 random trials and calculate the standard deviation of fitted values for $\bar{n}$ and $\Gamma$, which are obtained by averaging the results over either 20 or 200 trials.
-In our experiments, due to the control software’s delay in compiling a given set of control parameters, we choose to use 20 random trials, each trial repeated 100 times to obtain the qubit population.
-The experimental data indeed sit within the standard deviation for 20 trials.
-In future experiments with faster compilation Dalvi et al.
-(2023), we expect to perform a larger number of random trials with fewer repetitions each sup .
-With the capability of simulating a bath with continuous spectral density, we use multiple oscillators to simulate molecular energy transfer in a model structured bath.
-We simulate a bath of spectral density composed of up to 3 Lorentzian peaks, where each peak represents the vibrational mode of a molecular system such as light-harvesting protein complexes Olbrich et al.
-(2011); Chandrasekaran et al.
-(2015); Lee and Coker (2016); Kim et al.
-(2018); Tamascelli et al.
+Figure 1b illustrates the dynamics for various values of $\Gamma$ between 0 and $0.4\Delta$, which exhibit a phase transition from underdamped to overdamped dynamics as $\Gamma$ is increased (critical damping condition at $\Gamma\approx 0.18\Delta$).
+As the damping is increased beyond the critical point, the beat-note that signifies the excitation transfer from the spin to boson back to spin is suppressed, and there is a monotonic decrease in the oscillation of the donor population in the overdamped regime.
+We notice a $\pi$ phase flip in the oscillation of the donor population in the beat-note in the underdamped case (see, for example, the phase difference between the underdamped and overdamped case at $t\approx 2\pi/\Delta\times 10$).
+These features match very well with the theoretical predictions of spin-boson models.
+The uncertainty of $\bar{n}$ and $\Gamma$ realized in the experiment can be reduced by performing a larger number of random trials, each trial using a different set of random parameters.
+The shaded areas in Fig.
+1d and e show the standard deviation of fitted values for $\bar{n}$ and $\Gamma$, obtained by averaging the results over either 20 (pink) or 200 (blue) random trials.
+Our experiments are limited to 20 random trials due to the compiling delay required for each set of control parameters, and each trial is repeated 100 times to obtain the expectation value for the donor population.
+The experimental data is consistent with the standard deviation for 20 trials.
+We expect to readily increase the number of random trials with fewer repetitions each in future experiments with faster compilation tools Dalvi et al.
+(2023); sup .
+Engineering bath spectral densities
+We can use multiple oscillator modes to engineer a target spectral density structure for the oscillator bath.
+We simulate a bath of spectral density composed of up to 3 Lorentzian peaks, each represented by a motional mode of the ion chain.
+The spin-boson model with $J(\omega)$ in Fig.
+2b is simulated using the SDK operations over three motional modes.
+Fig.
+2a shows the frequency of the laser pulses used to drive the SDK, defined as the detuning of laser beams from the carrier transition frequency.
+The 1st, 3rd, and 5th motional modes are used, and the detuning of the laser frequency from the motional mode determines the frequency at which the bath spectral density is considered in Figure 2b.
+Here, we set the electronic coupling strength $\Delta=500$ cm-1 sup ; Chenu and Scholes (2015), such that at temperature 77 K, $\bar{n}\approx 0.1$ for the near-resonance vibrational modes.
+Figure 2c shows the results for simulating spin-boson models with spectral densities composed of 2 and 3 Lorentzian peaks, respectively.
+The theoretical predictions are obtained using the time-dependent density-matrix renormalization group algorithm White and Feiguin (2004); Prior et al.
+(2010) in the interaction picture Nuomin et al.
 (2022).
-We use the electronic coupling strength $\Delta=500$ cm-1 Chenu and Scholes (2015), such that at temperature 77 K, $\bar{n}\approx 0.1$ for the near-resonance vibrational modes.
+The donor population features underdamped coherent oscillations, with contributions from all three modes.
+The measured population closely follows the theoretical predictions.
+The coherence time of the motional modes in our setup of $\sim 5.18(67)$ ms is the leading source of error Wang et al.
+(2020); Kang et al.
+(2023a), and is comparable to the experimental time for simulating the time evolution of 2-mode (1.63 ms) and 3-mode (2.56 ms) cases.
+For these simulations, the contribution of the motional decoherence in the experiment is identical to the random kicks we apply in the spin-oscillator model to simulate the bath, and can be incorporated as the source of dephasing in our simulation with proper calibration Kang et al.
+(2023a); Milne et al.
+(2021); Keller et al.
+(2021).
 Figure 2: Simulations of molecular energy transfer in model structured baths.
@@
 The solid (dot-dashed) curves show the evolution with spectral densities given by the solid (dot-dashed) curves in (a).
-The spin-boson model with $J(\omega)$ in Fig.
-2b is simulated using the SDK operations of pulse frequencies shown in Fig.
-2a, where the pulse frequency is defined as the motion detuning of laser beams from the carrier transition.
-The 1st, 3rd, and 5th motional modes (in the order of increasing mode frequency) are used.
-Figure 2c shows the results for simulating spin-boson models with spectral densities composed of 2 and 3 Lorentzian peaks, respectively.
-The theoretical predictions are obtained using the time-dependent density-matrix renormalization group algorithm White and Feiguin (2004); Prior et al.
-(2010) in the interaction picture Nuomin et al.
-(2022).
-The oscillations in Fig.
-2c display repeated suppression and revival in amplitude, where the revival amplitude decreases with time, which is a signature of underdamped oscillations.
-The second suppression occurs earlier in the 3-mode curve than in the 2-mode curve, as the added Lorentzian peak increases the overall strength of spin-bath coupling.
-The suppressed oscillations are faster than before or after the suppression, as can be seen from the opposite oscillation phase between the 2-mode and 3-mode curves near $t\approx 650$ fs.
-Interestingly, the 3-mode experimental data exhibit slower oscillations during the second suppression.
-Nonetheless, with this exception, the experimental data overall closely align with the theoretical predictions, capturing the fine details of the oscillation envelope.
-The experimental time for simulating the time evolution is 1.63 ms (2.56 ms) for the 2-mode (3-mode) experiment, which is in the same order of magnitude as the motional coherence time $5.18(67)$ ms that characterizes motional dephasing, the leading source of errors in our trapped-ion system Kang et al.
-(2023a).
-Note that motional dephasing, when modelled as stochastic fluctuations of motional-mode frequencies Wang et al.
-(2020), simply increases the simulated value of $\Gamma$.
-A simulation of coherent spin-oscillator model with the same experimental time would have substantial errors due to motional dephasing Kang et al.
-(2024).
-However, for simulations of dephased spin-oscillator models, motional dephasing may have negligible effects, as the increase in $\Gamma$ may be small relative to the target value.
-To improve the accuracy of simulations with longer experimental time, the standard deviation of random frequency offsets may be reduced such that the effects of motional dephasing are compensated.
-This extends the capability of noisy hardware for simulating open quantum systems by essentially using the noise Guimarães et al.
-(2023); Leppäkangas et al.
-(2023).
-Reliable utilization of noise may require careful characterization, such as spectroscopy of motional dephasing Kang et al.
-(2023a); Milne et al.
-(2021); Keller et al.
-(2021).
-We further simulate the spin-boson models inspired by Leggett et al. Leggett et al. (1987). The spectral density is given by
+Simulating the Leggett spin-boson models
+We further simulate the spin-boson models inspired by Leggett et al. Leggett et al. (1987). The spectral density of the bath is given by
 $J_{\rm Legg}(\omega)=A\omega^{s}\omega_{c}^{1-s}e^{-\omega/\omega_{c}},$ (4)
-where $s<1$, $s=1$, and $s>1$ represent the sub-Ohmic, Ohmic and super-Ohmic baths, respectively, and $\omega_{c}$ is the cutoff frequency.
-This model is widely used to characterize noise in nanomechanical devices Seoanez et al.
+where $\omega_{c}$ is the cutoff frequency, and $s<1$, $s=1$, and $s>1$ represent the sub-Ohmic, Ohmic and super-Ohmic baths, respectively.
+This model is widely used to characterize the noise in nanomechanical devices Seoanez et al.
 (2007); Galland et al.
@@
 (2007).
-The spin dynamics are expected to exhibit a phase transition from coherent oscillations to incoherent decay or localization as $s$ is decreased or $A$ is increased, which is a subject of longstanding research using analytical tools Leggett et al.
+The spin dynamics are expected to exhibit a phase transition from coherent oscillations to incoherent decay or localization as $s$ is decreased or $A$ is increased, which has been a subject of longstanding research using analytical tools Leggett et al.
 (1987); Bulla et al.
@@
 (2023).
-Adding quantum simulators to the list of methods for studying this model is an interesting task.
-We use a sum of Lorentzian spectral densities Meier and Tannor (1999) to approximate the spectral density $J_{\rm Legg}(\omega)$.
+We use a sum of Lorentzian lines from motional modes to approximate the spectral density $J_{\rm Legg}(\omega)$ Meier and Tannor (1999).
 To find the optimal set of parameters $\kappa$’s, $\Gamma$’s, and $\nu$’s in Eq.
-(3) that best represent $J_{\rm Legg}(\omega)$, well-established global optimization algorithms Wolz (2022) are used to minimize the cost function, which is defined as the integral of the absolute difference between $J_{\rm Legg}(\omega)$ and the sum of corresponding Lorentzian spectral densities over a given bandwidth sup .
-As a first step toward simulating Leggett’s spin-boson models, we consider weak spin-bath coupling ($A=0.1$) and minimize the cost function for a limited bandwidth near resonance $\omega\in[0.9\Delta,1.1\Delta]$ using 2 Lorentzian peaks.
-Figure 3a shows the obtained sums of Lorentzian peaks (solid lines) that approximate sub-Ohmic ($s=0.5$), Ohmic ($s=1.0$), and super-Ohmic ($s=2.0$) spectral densities (dot-dashed lines), respectively.
-The results for simulating the spin-boson models with these approximated spectral densities are shown in Fig.
+(3) that best represent $J_{\rm Legg}(\omega)$, well-established global optimization algorithms Wolz (2022) are used to minimize the cost function, defined as the integral of the absolute difference between $J_{\rm Legg}(\omega)$ and the sum of corresponding Lorentzian spectral densities over a given bandwidth sup .
+As a first step, we consider weak spin-bath coupling ($A=0.1$) and designed the bath spectrum for a limited bandwidth near resonance $\omega\in[0.9\Delta,1.1\Delta]$ using 2 Lorentzian peaks.
+Figure 3a shows the obtained spectra from the modes (solid lines) that approximate sub-Ohmic ($s=0.5$), Ohmic ($s=1.0$), and super-Ohmic ($s=2.0$) spectral densities (dot-dashed lines), respectively.
+The results for the donor population dynamics interacting with a bath featuring these approximated spectral densities are shown in Fig.
 3b-d.
-We observe a crossover from near-coherent ($s=2$) to fully damped ($s=0.5$) oscillations in the simulated time scale.
+We observe a crossover from near-coherent ($s=2$) to fully damped ($s=0.5$) oscillations in the simulated time scale depending on the level of noise in the bath.
 The experimental data match well with theoretical predictions, where the deviations in the $s=2$ case result from insufficient number of Trotterization steps sup .
@@
 (b) Comparison of the population dynamics generated by the sub-Ohmic spectral density (dot-dashed blue) and its 10-mode approximation (solid red).
-Figure 3e compares the predicted population dynamics of the spin-boson model with spectral density $J_{\rm Legg}(\omega)$ and the corresponding approximated spectral density.
-While the dynamics match very well for the Ohmic and super-Ohmic cases, the sub-Ohmic case shows a non-negligible deviation.
-This is because the $J(\omega)$ values are overall larger for the sub-Ohmic bath, so matching the population dynamics requires accurately capturing a larger bandwidth.
-Indeed, in future experiments with a larger number of motional modes, we expect to match a wider range of spectral density and thus more accurately simulate the sub-Ohmic spin-boson model, as suggested by the 10-mode example in Fig.
+Figure 3e compares the predicted population dynamics of the spin-boson model with ideal spectral density $J_{\rm Legg}(\omega)$ and the corresponding approximated spectral density.
+While the dynamics match very well for the Ohmic and super-Ohmic cases, the sub-Ohmic case shows a non-negligible deviation due to the gap in the two spectral densities within the bandwidth (insets in Fig.
+3a).
+In future experiments with a larger number of motional modes Porras et al.
+(2008), we expect to match a wider range of spectral densities and thus more accurately simulate the sub-Ohmic spin-boson model, as suggested by the 10-mode numerical example in Fig.
 4.
-Note that Ref.
-Porras et al.
-(2008) suggests using roughly 100 motional modes to represent the Ohmic spectral density.
-Using noisy modes allows simulating the spin-boson models much more efficiently in terms of the number of motional modes used.
+Vibration-assisted energy transfer
 Finally, we simulate the VAET model Gorman et al.
 (2018) with nonzero energy detuning $\epsilon$ between the spin states.
+Here we set the spin parameters as $\Delta=30\>\text{cm}^{-1}$ and $\epsilon=100\>\text{cm}^{-1}$.
 In the absence of coupling to a vibrational mode ($\kappa=0\>\text{cm}^{-1}$), energy transfer between the detuned spin states is suppressed; however, for nonzero $\kappa$, the energy transfer can be activated if the mode frequency $\nu$ meets the resonance condition.
-Here we set the spin parameters as $\Delta=30\>\text{cm}^{-1}$ and $\epsilon=100\>\text{cm}^{-1}$.
 Figure 5: Simulations of vibrationally-assisted energy transfer (VAET).
-$\epsilon=100\>\text{cm}^{-1}$ and $\Delta=30\>\text{cm}^{-1}$ are fixed.
+We set $\epsilon=100\>\text{cm}^{-1}$ and $\Delta=30\>\text{cm}^{-1}$.
 (a) (b) Time evolution of the donor-state population for coherent VAET models with $\kappa=30\>\text{cm}^{-1}$.
@@
 Circles represent experimental data and error bars are standard deviation over 100 repetitions.
-In (a), bath temperature is fixed to 0 K ($\bar{n}=0$) and $\nu$ is varied.
+In (a), bath temperature is fixed at 0 K ($\bar{n}=0$) and $\nu$ is varied.
 In (b), $\nu$ is fixed to the resonant value $104\>\text{cm}^{-1}$ and two temperatures 0 K and 300 K, which correspond to $\bar{n}=0$ and 2.0, respectively, are simulated.
@@
 For $\kappa=10$ and $20\>\text{cm}^{-1}$, error bars indicate the standard deviation over 20 trials, each trial performed with randomly drawn set of parameters and repeated 100 times.
-We first explore the coherent spin-oscillator model $\hat{H}_{SO}$ with a relatively large coupling strength $\kappa=30\>\text{cm}^{-1}$ and various values of $\nu$.
-When $\nu$ satisfies the resonance condition of VAET, i.e., $\nu=\sqrt{\Delta^{2}+\epsilon^{2}}/k\approx 104\>\text{cm}^{-1}/k$ where $k$ is an integer, energy transfer occurs Gorman et al.
+Figure 5a shows the experimental data for simulating the coherent spin-oscillator model $\hat{H}_{SO}$ with a relatively large coupling strength ($\kappa=30\>\text{cm}^{-1}$).
+When $\nu$ satisfies the resonance condition of VAET ($\nu=\sqrt{\Delta^{2}+\epsilon^{2}}/k\approx 104\>\text{cm}^{-1}/k$ where $k$ is an integer), energy transfer occurs Gorman et al.
 (2018).
 Otherwise, energy remains localized at the donor site.
-Figure 5a shows the experimental data for simulating this model for the resonant ($\nu=52$ and $104\>\text{cm}^{-1}$) and non-resonant ($\nu=78\>\text{cm}^{-1}$) cases.
-Furthermore, in Fig.
-5b, we simulate the impact of initial temperature on the resonant ($\nu=104\>\text{cm}^{-1}$) VAET.
-We simulate baths of two temperatures 0 K and 300 K, which correspond to initial average phonon number $\bar{n}=0$ and $2.0$, respectively.
-For the latter case, SDK operations with random phase are applied to prepare the thermal state (when averaged over many trials).
-At first, the energy transfer is faster for the case of higher temperature, as more vibrational quanta assist the energy transfer.
-However, at later times, high temperature leads to a rapid decay of oscillations, as a thermal state of large $\bar{n}$ is a highly mixed state which quickly loses the coherent behavior.
-This pattern is accurately captured by our trapped-ion simulations.
-Noise (dephasing or damping) can be added to the VAET model.
+Figure 5b, shows the impact of initial temperature of the oscillators on the resonant VAET ($\nu=104\>\text{cm}^{-1}$).
+We simulate baths of two temperatures 0 K and 300 K corresponding to initial average phonon number $\bar{n}=0$ and $2.0$, respectively (see Methods).
+The energy transfer is faster for higher temperature bath in the beginning, as more vibrational quanta assist the energy transfer.
+However, high temperature leads to a rapid decay of oscillations over time as the coherence is lost.
 In Fig.
-5c we simulate the dephased spin-oscillator model with $\Gamma=10\>\text{cm}^{-1}$ and various values of $\kappa$, showing good match with theoretical predictions.
-However, as $\kappa$ is increased, the results deviate more from the theoretical predictions of the damped spin-oscillator model, which is non-perturbatively equivalent to those of the spin-boson model with $J(\omega)=J_{\rm Lo}(\omega)$ Lemmer et al.
+5c, we simulate the dephased spin-oscillator model with $\Gamma=10\>\text{cm}^{-1}$ and various values of $\kappa$, to study the impact of dissipative coupling.
+We note that as $\kappa$ is increased, the results deviate more from the theoretical predictions of the damped spin-oscillator model, which is non-perturbatively equivalent to those of the spin-boson model with $J(\omega)=J_{\rm Lo}(\omega)$ Lemmer et al.
 (2018); Tamascelli et al.
 (2018).
-This aligns with the aforementioned expectation that our method of using the dephased spin-oscillator model to approximate the spin-boson model is valid only in the weak-$\kappa$ regime.
-To summarize, the short-time dynamics between the two models match as the 2-point correlation functions are the same; however, the long-time dynamics deviate due to mismatch in the higher-order correlation functions sup .
-Indeed, for nonzero $\epsilon$, the steady-state donor populations are different between the dephased spin-oscillator model and the spin-boson model, as clearly shown by the $\kappa=20\>\text{cm}^{-1}$ curves.
-Discussion and outlook
-An immediate future direction is to observe the effects of noisy modes on the signature behaviors of small-scale molecular models using trapped-ion simulations Kang et al.
+This is consistent with the expectation that the method of using the dephased spin-oscillator model to approximate the spin-boson model is valid only in the weak-$\kappa$ regime.
+While the short-time dynamics between the two models match as the 2-point correlation functions are the same, the long-time dynamics deviate due to mismatch in the higher-order correlation functions for higher coupling strengths sup .
+Discussions and outlook
+Simulations of large-scale spin-boson models with strong coupling in the presence of dissipation is crucial yet challenging problem in physical Caldeira and Leggett (1983); Weiss et al.
+(1987); Schindler et al.
+(2022), chemical Garg et al.
+(1985); Gilmore and McKenzie (2005), and biological Huelga and Plenio (2013); Jang and Mennucci (2018) systems.
+While classical simulation approaches continue to improve with advancements in numerical methods, they encounter fundamental challenges when dealing with large quantum systems that span exponentially large Hilbert space MacDonell et al.
+(2021); Kang et al.
 (2024).
-Examples are: conical intersections Whitlow et al.
+Take, for example, the Dissipation-Assisted Matrix Product Factorization (DAMPF) method, a state-of-the-art method for simulating open quantum systems Somoza et al.
+(2019).
+For simulating a one-dimensional chain of $N_{s}$ electronic states each coupled to a bath of $N_{b}$ damped bosonic modes, the computational complexity of DAMPF scales as $\mathcal{O}(N_{s}^{5}N_{b}D_{b}^{2}\chi^{3})$, where $D_{b}$ is the number of energy levels that represent each bosonic mode and $\chi$ is the bond dimension in the matrix product operator representation of the system-bath density matrix.
+A critical issue is the rapid, potentially exponential increase in $\chi$ with respect to the system-bath coupling strength, for a fixed target accuracy in the population dynamics.
+For these types of problems, trapped-ion simulators may offer a more scalable alternative, as the number of operations scales only as $\mathcal{O}(N_{s}N_{b})$ and the total duration of SDK operations grows linearly with the system-bath coupling strength Kang et al.
+(2024).
+The current state-of-the-art trapped ion systems offer several dozens of highly coherent spins and bosonic modes that are very densely coupled with each other, the complete quantum description of which is totally intractable with classical computers of today Cetina et al.
+(2022); Chen et al.
+(2023b).
+Developing various control methods over these quantum systems to study important physical, chemical, biological and material systems can open up a new computational framework in science.
+The methods developed in this work utilizes fully programmable control methods to simulate dissipative processes.
+This approach can readily be applied to existing simulations of chemical processes to include coupling to environment Whitlow et al.
 (2023); Valahu et al.
-(2023), polarized-light-induced electron transfer Sun et al.
+(2023); Sun et al.
 (2023); Zhang et al.
-(2024), and VAET in the donor-bridge-acceptor system Li et al.
+(2024); Li et al.
 (2021).
-This may involve performing operations on a qudit Sun et al.
-(2023) or more than one qubits.
-For the latter case, entangling operations between the qubits Mølmer and Sørensen (1999) need to be added to the Trotterized instruction of operations.
-Our method of using randomized operations to simulate spin-boson models is limited to weak-coupling regime.
-However, simulating a bath of dephased oscillators may be of independent interest, as this can be thought as a natural extension from Gaussian ($\hat{L}\propto\hat{b}$ or $\hat{b}^{\dagger}$) to non-Gaussian ($\hat{L}\propto\hat{b}^{\dagger}\hat{b}$) bath models Linowski et al.
-(2022); Barthel and Zhang (2022).
-In such case, strong-coupling regime may be readily explored.
-For example, an interesting question is whether the phase transition from coherent oscillations to localization for sub-Ohmic spin-boson models Nalbach and Thorwart (2013); Duan et al.
+The dephased oscillator method can be used to realize non-Gaussian bath models Linowski et al.
+(2022); Barthel and Zhang (2022), relevant for studying physical processes such as phase transitions in strongly-coupled spin-boson models Duan et al.
 (2017); Otterpohl et al.
-(2022) can be observed using dephased spin-oscillator models that match the spectral densities.
-As evidenced in Fig.
-4, this may be simulated by our method applied to a few tens of motional modes, which can be readily utilized in a chain of 10 to 20 ions.
-Simulations of spin-boson models in strong-coupling regime can be performed using the damped spin-oscillator model, as its spin dynamics are non-perturbatively equivalent to those of the spin-boson model with the same spectral density Lemmer et al.
-(2018); Tamascelli et al.
-(2018).
-Damped oscillators can be realized using sympathetic cooling of the trapped ions’ motional modes Lemmer et al.
-(2018); Schlawin et al.
-(2021).
-This requires using a multi-species ion chain, such that the optical pumping during sympathetic cooling does not destroy the qubit states that encode the spins.
-If accomplished, this implementation of damping can be combined with our dephasing method to simulate unexplored models of open quantum systems.
-We note that while the role of electronic states’ dephasing in molecular energy transfer has been heavily studied Grad et al.
+(2022) and role of bosonic dephasing in molecular energy transfer Grad et al.
 (1988); Goldsmith et al.
@@
 (2010); Moreira et al.
-(2020), the role of vibrational modes’ dephasing is relatively less explored Yoneda et al.
+(2020); Yoneda et al.
 (2021).
-An ultimate goal is to accurately simulate the quantum dynamics of various models in physical Caldeira and Leggett (1983); Weiss et al.
-(1987), chemical Garg et al.
-(1985); Gilmore and McKenzie (2005), and biological Huelga and Plenio (2013); Jang and Mennucci (2018) systems that are intractable using classical simulation methods MacDonell et al.
-(2021); Kang et al.
-(2024).
-Classical simulation approaches continue to improve with advancements in numerical methods, yet they encounter fundamental challenges when dealing with large quantum systems.
-These challenges stem from the exponentially large Hilbert space and the quantum entanglement between the components.
-Take, for example, the Dissipation-Assisted Matrix Product Factorization (DAMPF) method Somoza et al.
-(2019), a state-of-the-art method for simulating open quantum systems.
-For simulating a one-dimensional chain of $N_{s}$ electronic states, each coupled to a bath of $N_{b}$ damped bosonic modes, the computational complexity of DAMPF scales as $\mathcal{O}(N_{s}^{5}N_{b}D_{b}^{2}\chi^{3})$, where $D_{b}$ is the number of energy levels that represent each bosonic mode and $\chi$ is the bond dimension in the matrix product operator representation of the system-bath density matrix Somoza et al.
-(2019).
-A critical issue is the rapid, potentially exponential increase in $\chi$ with respect to the system-bath coupling strength, for a fixed target accuracy in the population dynamics.
-In comparison, trapped-ion simulators may offer a more scalable alternative, as the number of operations scales as only $\mathcal{O}(N_{s}N_{b})$ and the total duration of SDK operations grows linearly with the system-bath coupling strength Kang et al.
-(2024).
-Also note that classical simulation methods typically struggle with complicated connectivity between the components, while for trapped ions all-to-all connectivity is granted both between qubits and between qubits and motional modes.
-Realizing accurate trapped-ion simulations of models with many spins and noisy bosonic modes requires overcoming more challenges, such as efficiently characterizing the motional-mode parameters Kang et al.
-(2023b) and significantly lowering the hardware’s noise level.
-Nonetheless, our demonstration of accurate simulations using controlled dissipation represents a key step forward in addressing these challenges.
-In particular, the accuracy of our results is enabled by carefully designed calibration schemes and precise phase tracking sup ; Lobser et al.
-(2023).
-Also, our method enables using motional dephasing noise for simulating open quantum systems, which significantly relaxes the limitation on the experiment time (for simulating the time evolution) of our simulations imposed by hardware noise.
-We anticipate this work inspires future works on improving the scale and accuracy of programmable quantum simulations of open quantum systems.
-Furthermore, our approach of finding models and parameter regimes that allow utilizing the dominant hardware noise may be of interest for various other analog quantum simulation platforms.
-During the preparation of our manuscript we became aware of Ref.
-So et al.
+High accuracy simulations will require development of efficient characterization methods Kang et al.
+(2023b) and novel control protocols in the trapped ion system sup ; Lobser et al.
+(2023) to implement various interaction terms in the target Hamiltonian.
+Combined with damped spin-boson models that can be implemented using sympathetic cooling Lemmer et al.
+(2018); Schlawin et al.
+(2021), the long-chain trapped ion systems can provide a versatile platform for simulating dynamics of complex open quantum systems.
+During the preparation of our manuscript we became aware of a complementary work So et al.
 (2024) which simulates the electron-transfer model using sympathetic cooling on the trapped ions’ motional mode.
-This complements our experiment on simulating spin-boson models with structured bath using randomized operations.
 Acknowledgements.
-The authors thank Yichao Yu for his valuable discussions about the experimental setup, Daniel Lobser for his valuable suggestion in controlling the RFSoC control software, and Thomas Barthel for his course on open quantum systems that inspired this work.
+The authors thank Yichao Yu for his valuable discussions about the phase tracking method, Daniel Lobser for his valuable suggestion in controlling the RFSoC control software, and Thomas Barthel for his course on open quantum systems that inspired this work.
 This research is funded by the Office of the Director of National Intelligence - Intelligence Advanced Research Projects Activity, through the ARO contract W911NF-16-1-0082 (provided the experimental apparatus used in the demonstration), the DOE BES award DE-SC0019400 (theoretical modeling), and the NSF Quantum Leap Challenge Institute for Robust Quantum Simulation Grant No.
-OMA-2120757 through a seed grant (experimental concepts and validation).
+OMA-2120757 (experimental concepts and validation).
 Support is also acknowledged from the U.S.
@@
 Figure 6: Experimental setup and instruction of operations.
-(a) Cartoon depiction of the experimental setup illustrates the orientation of components.
-The magnetic field is perpendicular to the trap surface, while the individual beam is perpendicular to the trap surface and the global beam is parallel to it.
-The differential wave vector $\Delta\textbf{k}$ is oriented at a 45-degree angle to the trap surface, aligning with one of the radial axes.
+(a) Schematic of the experimental setup illustrates the orientation of relevant fields.
+The magnetic field and the direction of the individual beam is perpendicular to the trap surface, while the global beam propagates is parallel to it.
+The differential wave vector $\Delta\textbf{k}$ is therefore oriented at a 45-degree angle to the trap surface, aligning with one of the radial axes.
 The global beam comprises two tones driving the blue (BSB) and red (RSB) sideband transitions, respectively.
 (b) Simplified energy levels of the 171Yb+ ion and the laser pulses used for implementing the on-resonance spin-dependent kick (SDK) operation.
-(c) Visualization of the motional-state evolution in phase space when a SDK operation of duration $\tau$ is applied, where the qubit is in the eigenstate of $\hat{\sigma}_{\phi_{s}}\equiv\hat{\sigma}_{+}e^{i\phi_{s}}+\hat{\sigma}_{-}e^{-i\phi_{s}}$.
+(c) Visualization of the motional-state evolution in phase space when a SDK operation of duration $\tau$ is applied, where the qubit is in the eigenstate of the operator $\hat{\sigma}_{\phi_{s}}\equiv\hat{\sigma}_{+}e^{i\phi_{s}}+\hat{\sigma}_{-}e^{-i\phi_{s}}$.
 Here, $\hat{\sigma}_{+}$ ($\hat{\sigma}_{-}$) is the raising (lowering) operator of the qubit, $\phi_{s}$ ($\phi_{m}$) is the spin (motion) phase of the laser beams, $\tilde{\Omega}$ is the Rabi frequency of the sideband transition, and $\delta_{m}$ is the motion detuning.
@@
 6a.
-The spin states are encoded to the 2S1/2-orbital hyperfine energy levels (qubit states) as |$\ket{0}\equiv\ket{F=1;m_{F}=0}$ and |$\ket{1}\equiv\ket{F=0;m_{F}=0}$.
+The spin states are encoded in the hyperfine energy levels in the 2S1/2 manifold (qubit states) as |$\ket{0}\equiv\ket{F=1;m_{F}=0}$ and |$\ket{1}\equiv\ket{F=0;m_{F}=0}$.
 The bosonic bath modes are encoded as the collective radial motional modes of the ion chain.
@@
 (2020).
-Acousto-optic modulators are employed to manipulate the frequency and phase of each laser beam, thereby enabling the execution of Raman transitions at specified frequency detunings and phases.
-Specifically, the spin-oscillator coupling is simulated using the spin-dependent kick (SDK) operation induced by simultaneous blue- and red-sideband transitions as shown in Fig.
+Acousto-optic modulators are employed to manipulate the frequency and phase of each laser beam, thereby enabling single-qubit gates: we apply a bit flip ($X$) gate to initialize the ion to |$\ket{0}$ state.
+The spin-oscillator coupling is simulated using the spin-dependent kick (SDK) operation induced by simultaneous blue- and red-sideband transitions as shown in Fig.
 6b.
-Figure 6c visualizes how the motional state evolves in phase space when a SDK operation of duration $\tau$, spin (motion) phase $\phi_{s}$ ($\phi_{m}$), motion detuning $\delta_{m}$ (see Fig.
-6b), and sideband Rabi frequency $\tilde{\Omega}$ is applied Lee et al.
+Figure 6c visualizes how the motional state evolves in phase space when a SDK operation of duration $\tau$, spin (motion) phase $\phi_{s}$ ($\phi_{m}$), motion detuning $\delta_{m}$, and sideband Rabi frequency $\tilde{\Omega}$ is applied Lee et al.
 (2005).
@@
 This method can be straightforwardly extended to the case where multiple modes are used; however, the laser phase needs to be tuned to track the phase of each motional mode correctly sup .
-The Trotterization method here enables simulating the VAET model on a single qubit.
-This is different from Ref.
-Gorman et al.
-(2018), which simulates the same model using continuous evolution of two ions induced by global laser beams.
-Thus, Trotterization using individually addressed beams is more efficient in terms of the required number of qubits.
-Also, our approach can be straightforwardly extended to simulating more complicated molecular models consisting of many electronic states and vibrational modes Kang et al.
+This approach can be readily extended to simulating more complicated molecular models consisting of many electronic states and vibrational modes Kang et al.
 (2024).
 Randomized operations
-Randomizing the control parameters is a key ingredient of our simulations of spin-boson models.
-By taking an average of many coherent evolutions, each involving a stochastically varying parameter, we realize both heating and dephasing as described by the Lindblad master equation Cai and Barthel (2013); Chenu et al.
+Randomizing the control parameters is a key ingredient in our approach to simulating dissipation mechanisms in spin-boson models.
+By taking an average of many coherent evolutions each involving a stochastically varying parameter, we realize both heating and dephasing as described by the Lindblad master equation Cai and Barthel (2013); Chenu et al.
 (2017).
 Heating (dephasing) is used to tune the temperature (spectral linewidth) of the bath.
-First, infinite-temperature heating, described by the pair of Lindblad operators $\hat{L}_{1}=\sqrt{\Gamma^{\prime}}\hat{b}$ and $\hat{L}_{2}=\sqrt{\Gamma^{\prime}}\hat{b}^{\dagger}$, is induced by applying SDK operations with stochastically varying motion phase $\phi_{m}$.
-Specifically, we apply $N^{\prime}$ resonant ($\delta_{m}=0$) SDK operations of duration $\tau^{\prime}=4\Gamma^{\prime}/\tilde{\Omega}^{2}$, spin phase $\phi_{s}=0$, and $\phi_{m}$ randomly drawn from $[0,2\pi)$ for each operation.
+First, the bosonic modes at finite temperature is created by modeling heating process where the modes interact with an infinite-temperature external bath for a finite duration, described by aa pair of Lindblad operators $\hat{L}_{1}=\sqrt{\Gamma^{\prime}}\hat{b}$ and $\hat{L}_{2}=\sqrt{\Gamma^{\prime}}\hat{b}^{\dagger}$.
+We apply $N^{\prime}$ resonant ($\delta_{m}=0$) SDK operations of duration $\tau^{\prime}=4\Gamma^{\prime}/\tilde{\Omega}^{2}$, spin phase $\phi_{s}=0$, and motion phase $\phi_{m}$ randomly drawn from $[0,2\pi)$ for each operation.
 These stochastic operations are applied on the initial qubit-mode composite state |$\ket{+,n=0}$; note that $\ket{+}$ is an eigenstate of the spin operator with $\phi_{s}=0$, and thus the qubit is decoupled from these operations.
 When averaged over many trials of coherent evolutions, each executed with a distinct set of $N^{\prime}$ random $\phi_{m}$ values, this procedure prepares the thermal state with an average phonon number $\bar{n}=N^{\prime}\Gamma^{\prime}\tau^{\prime}$.
-This can be intuitively understood as the thermal state represented as an ensemble of randomly displaced coherent states.
-See the Supplemental Material sup for derivations.
+This can be intuitively understood as the thermal state represented as an ensemble of randomly displaced coherent states (derivation provided in Supplementary Material sup ).
 For the experimental data in Fig.
@@
 5b, the average phonon number $\bar{n}$ of the initial thermal state is tuned by varying the duration $\tau^{\prime}$ of SDK operations, where the number of steps $N^{\prime}$ is fixed to 5.
-The errors in $\bar{n}$ due to finite number of steps is discussed in the Supplemental Material sup .
+The errors in $\bar{n}$ due to finite number of steps is discussed in the Supplementary Material sup .
 Next, dephasing during the time evolution, which is described by a single Lindblad operator $\hat{L}_{1}=\sqrt{\Gamma}\hat{b}$, is induced by giving random offsets to the motion detuning $\delta_{m}$ of the $N$ Trotterized SDK operations.
@@
 This enables tuning the FWHM of each Lorentzian peak in the bath’s spectral density.
-The errors in the simulated dephasing strength due to finite number of Trotterization steps is discussed in the Supplemental Material sup .
+The errors in the simulated dephasing strength due to finite number of Trotterization steps is discussed in the Supplementary Material sup .

### v3

@@
 Abstract
-The spin-boson model, involving spins interacting with a bath of quantum harmonic oscillators, is a widely used representation of open quantum systems.
-Trapped ions present a natural platform for simulating the quantum dynamics of such models, thanks to the presence of both high quality internal qubit states and the motional modes of the ions that can simulate the relevant quantum degrees of freedom.
-In our work, we extend the previous body of work that focused on coherent coupling of the spins and bosons to perform quantum simulations with structured dissipative baths using the motional states of trapped ions.
-We demonstrate the capability for adjusting the bath’s temperature and continuous spectral density by adding randomness to fully programmable control parameters.
-Subsequently, we simulate the dynamics of various spin-boson models with noise spectral densities constructed from coupling to several dissipative harmonic oscillator modes.
-The experimental outcomes closely align with theoretical predictions, indicating successful simulation of open quantum systems using a trapped-ion system.
-Interaction of a quantum system with its environment in an open dissipative system can significantly influence the dynamics of the system.
+The spin-boson model, involving spins interacting with a bath of quantum harmonic oscillators, is a widely used representation of open quantum systems that describe many dissipative processes in physical, chemical and biological systems.
+Trapped ions present an ideal platform for simulating the quantum dynamics of such models, by accessing both the high-quality internal qubit states and the motional modes of the ions for spins and bosons, respectively.
+We demonstrate a fully programmable method to simulate dissipative dynamics of spin-boson models using a chain of trapped ions, where the initial temperature and the spectral densities of the boson bath are engineered by controlling the state of the motional modes and their coupling with qubit states.
+Our method provides a versatile and precise experimental tool for studying open quantum systems.
+A new method to simulate open quantum systems made of spins and harmonic oscillators where the bath properties are engineered.
+Interaction of a quantum system with its environment in an open dissipative setting can significantly influence its dynamics.
 A paradigmatic model of non-Markovian open quantum systems that captures the quantum nature of both the system and the environment is the spin-boson model Leggett et al.
 (1987), which is ubiquitous in the study of condensed-matter physics Caldeira and Leggett (1983); Weiss et al.
-(1987); Schindler et al.
 (1987), chemical reactions Garg et al.
@@
 (2012), and biological light-harvesting complexes Huelga and Plenio (2013); Jang and Mennucci (2018).
-Experimentally realizing the dynamics of spin-boson models in a programmable fashion is an interesting challenge for quantum simulators Mostame et al.
+Experimentally realizing the dynamics of spin-boson models in a programmable fashion is an exciting challenge Mostame et al.
 (2012); Leppäkangas et al.
@@
 (2018); Kim et al.
-(2022).
-Trapped ions provide a highly controllable system of spins and motional modes, and thus are a natural platform for simulating spin-boson models.
+(2022), which can lead to new quantum simulation approaches that reach beyond classically tractable limit.
+Trapped ions provide a highly controllable system of spins and bosonic motional modes, and thus are a natural platform for simulating spin-boson models.
 The bosonic bath can be simulated using either a large number ($\gtrsim 50$) of motional modes Porras et al.
@@
 (2024).
-While significant experimental progress has been made in controlling the qubits and motional modes together Jia et al.
-(2022); Chen et al.
-(2023a), past works in this direction are limited to simulating a spin coherently coupled to up to two bosonic modes Gorman et al.
+While significant experimental progress has been made in controlling the qubits and motional modes together Chen et al.
+(2023a); Flühmann and Home (2020); Jia et al.
+(2022), past works in this direction are limited to simulating a spin coherently coupled to up to two bosonic modes Whitlow et al.
+(2023); Valahu et al.
+(2023); Gorman et al.
 (2018); MacDonell et al.
-(2023); Whitlow et al.
-(2023); Valahu et al.
 (2023).
-These previous studies did not consider the effects of dissipation, which is often described by the coupling of a system to a continuum of bath modes, that can determine the system’s long-term behavior such as reaching a steady state.
-In this work, we perform quantum simulations of spin-boson models with continuous spectral densities using trapped ions and their motional modes.
-Specifically, we simulate spectral densities composed of several Lorentzian peaks, which is characteristic of baths in light-harvesting protein complexes Olbrich et al.
-(2011); Chandrasekaran et al.
-(2015); Lee and Coker (2016); Kim et al.
-(2018); Tamascelli et al.
+These studies could not consider the effects of dissipation often described by coupling of the system to a continuum of bath modes, which determines the system’s long-term behavior such as reaching a steady state.
+Dissipative couplings come in two flavors: damping is when the quantum system exchanges energy with the bath (inelastic coupling) and eventually reaches thermal equilibrium with the bath, while dephasing is when the phase of the quantum system is randomized by interaction with the bath without energy exchange (elastic coupling).
+In this work, we perform quantum simulations of spin-boson models with programmable spectral densities using a 7-ion chain and up to 3 of its motional modes.
+The target spectral densities of the bath are decomposed into several Lorentzian peaks, which is characteristic of baths in real-world systems such as light-harvesting protein complexes Lee and Coker (2016); Kim et al.
 (2018).
-Furthermore, we simulate the sub-Ohmic, Ohmic, and super-Ohmic spin-boson models Leggett et al.
-(1987), which are ubiquitous in various open quantum systems ranging from nanomechanical devices Seoanez et al.
-(2007); Galland et al.
-(2008); Chen et al.
-(2011) and superconducting circuits Magazzù et al.
-(2018); Thorwart et al.
-(2004); Kaur et al.
-(2021); Gulácsi and Burkard (2023) to molecules in which proton-transfer reactions occur Cukier and Morillo (1989); Ohta et al.
-(2006); Craig et al.
+Our approach can readily simulate the dephasing model of dissipation, by adding randomness to the control parameters of the laser that drives the spin-boson interaction, such as frequency and phase.
+Using this technique, we study the dynamics of classic spin-boson models Leggett et al.
+(1987) and the vibration-assisted energy transfer (VAET) process Gorman et al.
 (2018).
-Finally, we simulate the vibrationally-assisted energy transfer (VAET) Gorman et al.
-(2018) in a simple molecular model where dephasing is added to the vibrational mode.
-The various spin-state population dynamics, which display behaviors such as damped oscillations, coherent revivals, and relaxation to a steady state, are accurately captured by our quantum-simulation results.
-We note that a recent study Wang et al.
-(2024) simulated a spin dissipatively coupled to many modes, albeit with limited tunability.
-Capturing the dissipative behavior of the spin-boson model, as well as tuning the bath’s temperature, is enabled by adding randomness to the control parameters, such as the laser frequency and phase.
-Remarkably, the impact of these control signals is directly analogous to motional dephasing, the leading source of noise in our experimental setup Kang et al.
-(2023a).
-This serves as a lower limit for the level of noise that can be introduced in our quantum simulation process.
-While the quantum simulations presented here are limited to the weak spin-bath coupling regime, our demonstration of adding or using noise Guimarães et al.
-(2023); Leppäkangas et al.
-(2023) in analog quantum simulations is an important ingredient for a potential quantum advantage in simulating open quantum system models that are intractable using classical simulation methods MacDonell et al.
+Our demonstration to engineer spectral density of the bath in analog quantum simulations is an important ingredient for achieving quantum advantage in simulating open quantum system models that are intractable using classical methods MacDonell et al.
 (2021); Kang et al.
 (2024).
-Results
-Models
+We note that a recent study Wang et al. (2024) simulated a spin dissipatively coupled to many modes, albeit with limited tunability.
+Dephased spin-oscillator model
 The spin-boson model describes a spin coupled to a continuous bath of quantum harmonic oscillators. The Hamiltonian is given by
@@
 […]
-A spin-boson model with a structured bath is often characterized by a spectral, density $J(\omega)$, composed of multiple Lorentzian peaks, each of which can be represented by a noisy harmonic oscillator Lemmer et al.
+We decompose the spectral density $J(\omega)$ of a structured bath into a sum of multiple Lorentzian peaks, each of which can be represented by a dissipative harmonic oscillator Lemmer et al.
 (2018).
-Specifically, we consider a spin coupled to dephased oscillators (hereafter referred to as the dephased spin-oscillator model).
-The single-oscillator case is described by the Hamiltonian
-$\hat{H}_{SO}=\hat{H}_{S}+\frac{\kappa}{2}\hat{\sigma}_{Z}\otimes(\hat{b}+\hat{b}^{\dagger})+\nu\hat{b}^{\dagger}\hat{b}$ (2)
-and a single Lindblad operator $\hat{L}_{1}=\sqrt{\Gamma}\hat{b}^{\dagger}\hat{b}$, where $\hat{b}$ is the annihilation operator of the oscillator, such that the composite state $\rho$ of spin and oscillators follows the Lindblad master equation $\dot{\rho}=-i[\hat{H}_{SO},\rho]+\sum_{j}(\hat{L}_{j}\rho\hat{L}_{j}^{\dagger}-\frac{1}{2}\{\hat{L}_{j}^{\dagger}\hat{L}_{j},\rho\})$ Lindblad (1976).
-In a reasonable regime of parameters ($\Gamma<\nu/2$, $\beta\ll 2\pi(\Gamma/2)^{-1}$) Lemmer et al.
+In this work, we consider a spin coupled to several oscillators subject to constant dephasing (“dephased” spin-oscillator model).
+This “discrete” oscillator model is described by the Hamiltonian
+$\hat{H}_{D}=\hat{H}_{S}+\sum_{l}\frac{\kappa_{l}}{2}\hat{\sigma}_{Z}\otimes(\hat{b}_{l}+\hat{b}_{l}^{\dagger})+\sum_{l}\nu_{l}\hat{b}_{l}^{\dagger}\hat{b}_{l}$ (2)
+and corresponding Lindblad operators $\hat{L}_{l}=\sqrt{\Gamma}_{l}\hat{b}_{l}^{\dagger}\hat{b}_{l}$, where $\kappa_{l}$ represents the coupling strength between the spin and $l$-th oscillator, $\nu_{l}$ and $\Gamma_{l}$ denote the frequency and the dephasing rate of the $l$-th oscillator, respectively, and $\hat{b}_{l}$ is the annihilation operator of the $l$-th oscillator.
+The composite state $\rho$ of spin and oscillators follows the Lindblad master equation $\dot{\rho}=-i[\hat{H}_{D},\rho]+\sum_{l}(\hat{L}_{l}\rho\hat{L}_{l}^{\dagger}-\frac{1}{2}\{\hat{L}_{l}^{\dagger}\hat{L}_{l},\rho\})$.
+In a reasonable regime of parameters ($\Gamma_{l}<\nu_{l}/2$, $\beta\ll 2\pi(\Gamma_{l}/2)^{-1}$) Lemmer et al.
 (2018), this bath of dephased oscillators is assigned a spectral density composed of Lorentzian peaks sup
-$J_{\rm Lo}(\omega)=\kappa^{2}\Big{(}\frac{\Gamma/2}{(\Gamma/2)^{2}+(\omega-\nu)^{2}}-\frac{\Gamma/2}{(\Gamma/2)^{2}+(\omega+\nu)^{2}}\Big{)},$ (3)
-where we assume that the oscillator is initially in the thermal state with average phonon number $\bar{n}(\nu)$.
-The spectral density of the bath determines the dynamics of the spin to leading order in $\kappa T$, where $T$ is the evolution time Breuer and Petruccione (2002); sup .
-Therefore, the spin dynamics of the dephased spin-oscillator model approximately match those of the spin-boson model with $J(\omega)=J_{\rm Lo}(\omega)$ in the weak-$\kappa$ regime, which extends straightforwardly to the model with multiple oscillators.
-Of particular interest is the unbiased ($\epsilon=0$) spin-boson model Leggett et al.
+$J_{\rm Lo}(\omega)=\sum_{l}\kappa_{l}^{2}\Big{(}\frac{\Gamma_{l}/2}{(\Gamma_{l}/2)^{2}+(\omega-\nu_{l})^{2}}-\frac{\Gamma_{l}/2}{(\Gamma_{l}/2)^{2}+(\omega+\nu_{l})^{2}}\Big{)},$ (3)
+where we assume that each oscillator is initially in the thermal state with average phonon number $\bar{n}(\nu_{l})$.
+The spectral density of the bath determines the dynamics of the spin to leading order in $\bar{\kappa}T$, where $\bar{\kappa}$ is the average coupling strength and $T$ is the evolution time sup .
+Therefore, the spin dynamics of the dephased spin-oscillator model approximately match those of the spin-boson model with $J(\omega)=J_{\rm Lo}(\omega)$ in the weak-$\bar{\kappa}$ regime.
+In the unbiased ($\epsilon=0$) spin-boson model Leggett et al.
 (1987); Wang and Thoss (2008); Duan et al.
 (2017); Strathearn et al.
-(2018), where the equilibrium donor-state population is $1/2$ for both dephased spin-oscillator models and spin-boson models.
-Thus, for weak spin-bath coupling, the population dynamics of the two models show an excellent match at all times; the match in short (long) time scale is achieved due to matching spectral densities (equilibrium populations).
-Unless specified otherwise, we use $\epsilon=0$; results for $\epsilon\neq 0$ can be found in Fig.
-5.
-A more commonly considered model is a spin coupled to damped oscillators (hereafter referred to as the damped spin-oscillator model); for example, the single-oscillator case is described by the Hamiltonian $\hat{H}_{SO}$ and a pair of Lindblad operators $\hat{L}_{1}=\sqrt{\Gamma(\bar{n}(\nu)+1)}\hat{b}$ and $\hat{L}_{2}=\sqrt{\Gamma\bar{n}(\nu)}\hat{b}^{\dagger}$ Lemmer et al.
+(2018), the equilibrium donor-state population is $1/2$ for both models, and we expect the population dynamics of the two models show an excellent match at all times for weak spin-bath coupling.
+For the case of biased spin-boson models in the stronger coupling regime ($\epsilon\neq 0$, such as discussed in the VAET case), the dynamics of the dephased spin-oscillator model we simulate can deviate significantly from the traditional spin-boson model.
+Figure 1: Schematic of the experimental setup.
+The ions are confined in a micro-fabricated surface trap.
+The spin states are encoded as the hyperfine clock states (qubit states) and the bosonic bath modes are encoded as the collective radial motional modes of the ion chain.
+We strategically use the center ion and the three symmetric modes (except for the center-of-mass mode, as shown in the upper left corner) in the experiment to minimize cross-mode couplings sup .
+The laser frequencies for the SDK operations are determined by the frequency difference between the two Raman beams, and are depicted by the red and blue lines on the frequency axis detuned by $\delta_{m}$ from the mode frequency $\omega_{m}$, with the color gradients indicating the random frequencies used to implement the decoherence simulation.
+A more common model for dissipation is a spin coupled to oscillators subject to constant damping (“damped” spin-oscillator model) described by a pair of Lindblad operators $\hat{L}_{l1}=\sqrt{\Gamma_{l}(\bar{n}(\nu_{l})+1)}\hat{b}_{l}$ and $\hat{L}_{l2}=\sqrt{\Gamma_{l}\bar{n}(\nu_{l})}\hat{b}_{l}^{\dagger}$ Lemmer et al.
 (2018); Schlawin et al.
-(2021).
-For this model, a stronger result holds: the spin dynamics are non-perturbatively equivalent to those of the spin-boson model with the same spectral density, up to all orders in $\kappa T$ Tamascelli et al.
+(2021), where $\Gamma_{l}$ here is the damping rate of the $l$-th oscillator.
+A stronger result holds for this model: the spin dynamics are non-perturbatively equivalent to those of the spin-boson model with the same spectral density, up to all orders in $\bar{\kappa}T$ Tamascelli et al.
 (2018); sup .
@@
 (2021).
-In this work, we focus on dephased oscillators to represent the structured bath, which is readily performed using programmable randomized control operations on a single species of trapped ions.
-Single oscillator model
-The dephased spin-oscillator model can be simulated using trapped ions by applying spin-dependent kick (SDK) operations, which induce the coupling between the trapped-ion qubit and the motional mode (see Methods for details).
-By adding randomness to the control parameters of the SDK operations, we can also prepare the thermal state and simulate the dephasing of the motional mode, by averaging the results over many random trials.
-The key idea is that noisy (heated or dephased) evolution, described by the Lindblad master equation, is equivalent to an average of many coherent evolutions, each subject to a stochastic Hamiltonian Cai and Barthel (2013); Chenu et al.
-(2017).
-Figure 1: Demonstration of the tunability of the bath’s temperature and spectral linewidth.
-(a) Dynamics of the coherent spin-oscillator model $\hat{H}_{SO}$ with various values of average phonon number $\bar{n}$ of the oscillator’s initial state.
-(b) Dynamics of the spin-boson model with $\bar{n}=0$, $J(\omega)=J_{\rm Lo}(\omega)$ and various values of $\Gamma$.
-For both (a) and (b), $\epsilon=0$, $\nu=\Delta$, and $\kappa=0.1\Delta$.
-Lines (dots) represent theoretical predictions (experimental data) and error bars denote the standard deviation over 20 trials, each trial performed with randomly drawn set of parameters and repeated 100 times.
+The dephased oscillator model developed in this work expands the ability to simulate a broader set of bath models not experimentally accessible before.
+Experimental Implementation
+Figure 2: Demonstration of the tunability of the bath’s initial temperature and spectral linewidth.
+(a) Dynamics of the coherent spin-single oscillator model $\hat{H}_{D}$ with various values of average phonon number $\bar{n}$ for the oscillator’s initial state.
+(b) Dynamics of the model with $\bar{n}=0$, $J(\omega)=J_{\rm Lo}(\omega)$ (single peak) and various values of $\Gamma_{1}\equiv\Gamma$.
+For both (a) and (b), $\epsilon=0$, $\nu_{1}=\Delta$, and $\kappa_{1}=0.1\Delta$.
+Lines and dots represent theoretical predictions and experimental data, respectively.
+The error bars (size comparable to the symbols in these two plots) denote measured standard deviation over 20 trials, each trial performed with randomly drawn set of parameters and repeated 100 times.
 (c) Revival amplitude versus $\bar{n}$.
 Predictions of the original (solid line) model and the noise-aware (dashed line) model are compared with the experimental results (circles).
-(d) Fitted values of $\bar{n}$, where the experimental data are fitted to the predictions of the noise-aware model.
+(d) Measured values of $\bar{n}$ fitted to the predictions of the noise-aware model.
 Solid line depicts $y=x+\bar{n}_{0}$, where $\bar{n}_{0}=0.036$ is the expected offset in the phonon number due to imperfect cooling.
-(e) Fitted values of $\Gamma$, where the experimental data are fitted to the theoretical predictions with $\bar{n}=\bar{n}_{0}$.
+(e) Measured values of $\Gamma$ fitted to the theoretical predictions with $\bar{n}=\bar{n}_{0}$.
 Solid line depicts $y=x$.
-In (d) and (e), the blue (brown) shaded region represents the expected uncertainty of the respective value when the populations are averaged over 20 (200) random trials.
+In (d) and (e), the blue and brown shaded regions represent the expected uncertainty of the value when the populations are averaged over 20 and 200 random trials, respectively.
 (f) Lorentzian spectral densities for various target (solid) and fitted (dashed) values of $\Gamma$.
 The shaded region represents the expected uncertainty of $\Gamma$ when averaged over 20 random trials.
-We first demonstrate the programmability of the oscillator bath’s temperature, quantified by the average phonon number $\bar{n}$ of the bath oscillator’s initial state.
-Controlled amounts of resonant SDK operations with stochastically varying phase are applied to the ion in the motional ground state, which prepares the thermal state when averaged over many random trials (see Methods for details).
-Figure 1a illustrates the simulated coherent time evolution with respect to $\hat{H}_{SO}$, for various values of $\bar{n}$.
-The dynamics exhibit coherent oscillations with an envelope of slow beat-note that signifies resonant ($\nu=\Delta$) energy exchange between the system and the bath oscillator Eberly et al.
+The simulations of spin-boson models are performed using a linear chain of seven 171Yb+ ions (Figure 1).
+Two hyperfine internal states (|$\ket{0}$ and |$\ket{1}$) are used to represent the spin (or qubit).
+The ions sit 68 $\mu$m above the surface of the trap Revelle (2020), and the heating rate and decoherence time of the zig-zag motional mode are $3.6(3)$ quanta/s and $5.2(7)$ ms, respectively.
+The other motional modes used have similar magnitudes.
+Details of our experimental setup can be found in Ref.
+Wang et al.
 (2020).
-Our initial thermal state is prepared as an ensemble of randomly displaced coherent states.
+Using Doppler, electromagnetically-induced transparency (EIT), and sideband cooling techniques, the zig-zag motional mode can be cooled to near ground state with $\bar{n}=0.036(16)$.
+Standard qubit manipulation techniques can be used to initialize and measure the qubits, and single-qubit operations are driven by stimulated Raman transitions.
+The spin-oscillator coupling is simulated using the spin-dependent kick (SDK) operation induced by simultaneous application of blue- and red-sideband transitions sup .
+The evolution of $\hat{H}_{D}$ in Eq.
+(2) is mapped to a sequence of single-qubit rotations and SDK operations via Trotterization.
+We apply time evolution operator over a short time interval corresponding to each term in the Hamiltonian in interaction picture, and repeat them over many time intervals to simulate the evolution dynamics sup .
+This approach can be readily extended to simulating more complicated molecular models consisting of many electronic states and bath modes Sun et al.
+(2023); Kang et al.
+(2024).
+The dephased spin-oscillator model can be simulated using trapped ions by applying the SDK operations, which induce the coupling between the qubit and the motional mode.
+By adding randomness to the control parameters of the SDK operations and averaging the results over many random trials, we can implement dissipative processes such as preparing the thermal state and simulating the dephasing of the motional mode.
+The key idea is that certain dissipative evolutions described by the Lindblad master equation are equivalent to an average of many coherent evolutions, each subject to a stochastic Hamiltonian Chenu et al.
+(2017).
+Details of the procedure is described in Supplementary Material sup .
+Single-oscillator case
+We first consider a single-oscillator case ($l=1$), and demonstrate the programmability of the bath’s initial temperature quantified by the oscillator’s average phonon number $\bar{n}$.
+Controlled amounts of resonant SDK operations with stochastically varying phase are applied to the ion prepared near the motional ground state, which prepares the thermal state equivalent to an ensemble of randomly displaced coherent states sup .
+Figure 2a illustrates the simulated time evolution of the donor state population when it is coupled to such initial bath state, for various values of $\bar{n}$.
+The dynamics exhibit coherent oscillations with an envelope of slow collapse and revival (“beat-note”) that signifies resonant ($\nu_{1}=\Delta$) energy exchange between the system and the bath oscillator Eberly et al.
+(1980).
 A thermal state of larger $\bar{n}$ is a mixture of more “scattered” coherent states, leading to reduced beat-note amplitude at $t\approx 2\pi/\Delta\times 10.5$.
-We see a notable agreement between theoretical predictions (solid lines) and experimental results (dots).
 We compare our experimental results with the predictions of noise-aware models, reflecting the noise processes in the experimental setup.
-We use spin-boson models with initial phonon number $\bar{n}+\bar{n}_{0}$ and Lorentzian spectral density with full width at half maximum (FWHM) $\Gamma$ [see Eq.
-(3)], where the values $\bar{n}_{0}=0.036$ and $\Gamma=0.0022\Delta$ are extracted from separate experiments to characterize imperfect cooling and finite motional coherence time, respectively.
+We fit the experimental data using an oscillator model with initial phonon number $\bar{n}+\bar{n}_{0}$ and Lorentzian spectral density with full width at half maximum (FWHM) $\Gamma_{1}\equiv\Gamma$ [Eq.
+(3)], where the values $\bar{n}_{0}=0.036$ and $\Gamma=0.0022\Delta$ are extracted from independent measurements characterizing imperfect cooling and finite motional coherence time, respectively.
 The experimentally measured revival amplitude (Fig.
-1c) and fitted $\bar{n}$ (Fig.
-1d) match well with the predictions of noise-aware models.
-This shows that well-characterized experimental noise can be used as part of the dissipation model, and thus is not always a bottleneck for the accuracy of our simulations.
-Next, we show the tunability of the dephasing strength, which corresponds to the linewidth of the spin-boson model’s spectral density.
-Dephasing is implemented by adding random frequency offsets to the SDK operations that simulate the time evolution (see Methods for details).
+2c) and fitted $\bar{n}$ (Fig.
+2d) match well with the predictions of noise-aware models.
+This shows that well-characterized experimental noise can serve as the baseline values for the thermal excitation and dephasing rate in the open quantum system model, and thus is not always a bottleneck for the accuracy of our simulations.
+Next, we show the tunability of the dephasing rate $\Gamma$, which corresponds to the linewidth of the spin-boson model’s spectral density.
+Dephasing is implemented by adding random frequency offsets to the SDK operations that simulate the time evolution sup .
 This mimics the random fluctuations of the motional-mode frequency, following the stochastic model of motional dephasing noise in trapped-ion systems Wang et al.
 (2020).
-Figure 1b illustrates the dynamics for various values of $\Gamma$ between 0 and $0.4\Delta$, which exhibit a phase transition from underdamped to overdamped dynamics as $\Gamma$ is increased (critical damping condition at $\Gamma\approx 0.18\Delta$).
+Figure 2b illustrates the dynamics for various values of $\Gamma$ between 0 and $0.4\Delta$, which exhibit a crossover from underdamped to overdamped dynamics as $\Gamma$ is increased (critical damping condition at $\Gamma\approx 0.18\Delta$).
 As the damping is increased beyond the critical point, the beat-note that signifies the excitation transfer from the spin to boson back to spin is suppressed, and there is a monotonic decrease in the oscillation of the donor population in the overdamped regime.
-We notice a $\pi$ phase flip in the oscillation of the donor population in the beat-note in the underdamped case (see, for example, the phase difference between the underdamped and overdamped case at $t\approx 2\pi/\Delta\times 10$).
-These features match very well with the theoretical predictions of spin-boson models.
+We notice a $\pi$ phase flip in the oscillation of the donor population in the beat-note in the underdamped case, signified by the phase difference between the underdamped and overdamped cases at $t\approx 2\pi/\Delta\times 10$.
+Figure 3: Simulations of molecular energy transfer in model structured baths.
+(a) Simulated spectral density of the spin-boson model’s bath, where the electronic coupling strength $\Delta=500\>\text{cm}^{-1}$.
+The orange and brown solid curves illustrate the spectral densities summed over two and three peaks, respectively.
+(b) Dashed lines represent the radial motional modes that are coupled by the Raman beam.
+Black or grey dashed lines represent modes with non-zero or zero coupling strength to the target ion, respectively.
+The three solid lines and the shaded regions depict the mean value and the standard deviation of the randomized pulse frequencies, and the color corresponds to the Lorentzian peak generating the final spectral density in (a).
+(c) Time evolution of the spin-boson models with $J(\omega)$ in (a).
+The bath temperature is 77 K such that $\bar{n}\approx 0.1$, which is set for trapped ions’ motional modes using imperfect sideband cooling.
+Curves and dots represent theoretical predictions and experimental data, respectively.
+Error bars are derived as in Fig.
+2.
 The uncertainty of $\bar{n}$ and $\Gamma$ realized in the experiment can be reduced by performing a larger number of random trials, each trial using a different set of random parameters.
 The shaded areas in Fig.
-1d and e show the standard deviation of fitted values for $\bar{n}$ and $\Gamma$, obtained by averaging the results over either 20 (pink) or 200 (blue) random trials.
+2d and e show the standard deviation of fitted values for $\bar{n}$ and $\Gamma$, obtained by numerical simulation of results averaged over 20 (pink) or 200 (blue) random trials.
 Our experiments are limited to 20 random trials due to the compiling delay required for each set of control parameters, and each trial is repeated 100 times to obtain the expectation value for the donor population.
-The experimental data is consistent with the standard deviation for 20 trials.
+The experimental data is consistent with the standard deviation derived from 20 numerical trials.
 We expect to readily increase the number of random trials with fewer repetitions each in future experiments with faster compilation tools Dalvi et al.
@@
 Engineering bath spectral densities
-We can use multiple oscillator modes to engineer a target spectral density structure for the oscillator bath.
-We simulate a bath of spectral density composed of up to 3 Lorentzian peaks, each represented by a motional mode of the ion chain.
-The spin-boson model with $J(\omega)$ in Fig.
-2b is simulated using the SDK operations over three motional modes.
-Fig.
-2a shows the frequency of the laser pulses used to drive the SDK, defined as the detuning of laser beams from the carrier transition frequency.
-The 1st, 3rd, and 5th motional modes are used, and the detuning of the laser frequency from the motional mode determines the frequency at which the bath spectral density is considered in Figure 2b.
-Here, we set the electronic coupling strength $\Delta=500$ cm-1 sup ; Chenu and Scholes (2015), such that at temperature 77 K, $\bar{n}\approx 0.1$ for the near-resonance vibrational modes.
-Figure 2c shows the results for simulating spin-boson models with spectral densities composed of 2 and 3 Lorentzian peaks, respectively.
-The theoretical predictions are obtained using the time-dependent density-matrix renormalization group algorithm White and Feiguin (2004); Prior et al.
-(2010) in the interaction picture Nuomin et al.
-(2022).
-The donor population features underdamped coherent oscillations, with contributions from all three modes.
-The measured population closely follows the theoretical predictions.
-The coherence time of the motional modes in our setup of $\sim 5.18(67)$ ms is the leading source of error Wang et al.
-(2020); Kang et al.
-(2023a), and is comparable to the experimental time for simulating the time evolution of 2-mode (1.63 ms) and 3-mode (2.56 ms) cases.
-For these simulations, the contribution of the motional decoherence in the experiment is identical to the random kicks we apply in the spin-oscillator model to simulate the bath, and can be incorporated as the source of dephasing in our simulation with proper calibration Kang et al.
-(2023a); Milne et al.
-(2021); Keller et al.
-(2021).
-Figure 2: Simulations of molecular energy transfer in model structured baths.
-(a) Dashed lines represent the radial motional modes strongly coupled by the Raman beam.
-Black (grey) lines denote the modes with non-zero (zero) coupling strength to the target ion.
-The three solid lines (shaded regions) depict the mean value (standard deviation) of the randomized pulse frequencies.
-(b) Simulated spectral density of the spin-boson model’s bath, where the electronic coupling strength $\Delta=500\>\text{cm}^{-1}$.
-The dashed curves depict the Lorentzian peak generated by the pulses of corresponding colors in (a).
-The orange (brown) solid curves illustrate the spectral density summed over two (three) peaks.
-(c) Time evolution of the spin-boson models with $J(\omega)$ in (b).
-The bath temperature is 77 K such that $\bar{n}\approx 0.1$, which is set for trapped ions’ motional modes using imperfect sideband cooling.
-Curves (dots) represent theoretical predictions (experimental data).
-Error bars denote the standard deviation over 20 trials, each trial performed with randomly drawn set of parameters and repeated 100 times.
 Figure 4: Simulations based on Leggett’s spin-boson model.
@@
 Circles denote experimental data and error bars indicate the standard deviation over 60, 20, and 20 trials for (b), (c), and (d), respectively.
-Each trial is performed with randomly drawn set of parameters and repeated 100 times.
-(e) Time evolution of the spin-boson models.
-The solid (dot-dashed) curves show the evolution with spectral densities given by the solid (dot-dashed) curves in (a).
-Simulating the Leggett spin-boson models
+We can use multiple oscillator modes to engineer a target spectral density structure for the oscillator bath.
+We simulate a bath of spectral density composed of up to 3 Lorentzian peaks, each represented by a motional mode of the ion chain.
+The bath spectral density $J(\omega)$ in Fig.
+3a is simulated using the SDK operations over two or three motional modes.
+Figure 3b shows the frequencies of the laser pulses used to drive the SDK, defined as the detuning of laser beams from the carrier transition frequency.
+The 1st, 3rd, and 5th motional modes are used, and the detuning of the laser frequency from each motional mode determines the frequency at which the bath spectral density is considered in Fig.
+3a.
+We set the electronic coupling strength $\Delta=500$ cm-1 Chenu and Scholes (2015), such that at temperature 77 K, $\bar{n}\approx 0.1$ for the near-resonance vibrational modes.
+Figure 3c shows the results for simulating the dynamics of spin-boson models with spectral densities composed of 2 and 3 Lorentzian peaks, respectively.
+The theoretical predictions are obtained using the time-dependent density-matrix renormalization group algorithm in the interaction picture Nuomin et al.
+(2022).
+The donor population features underdamped coherent oscillations, with contributions from all three modes.
+The measured population closely follows the theoretical predictions.
+The coherence time of the motional modes in our setup of $\sim 5.2(7)$ ms is the leading source of noise Wang et al.
+(2020); Kang et al.
+(2023a), and is comparable to the experimental time for simulating the time evolution under spectral densities generated from 2 mode (1.63 ms) and 3-mode (2.56 ms) cases.
+For these simulations, the contribution of motional decoherence in the experiment is identical to the random kicks that we apply in the spin-oscillator model to simulate the bath, and can be incorporated as the source of dephasing in our simulation with proper calibration Kang et al.
+(2023a).
+Leggett spin-boson models
 We further simulate the spin-boson models inspired by Leggett et al. Leggett et al. (1987). The spectral density of the bath is given by
@@
 where $\omega_{c}$ is the cutoff frequency, and $s<1$, $s=1$, and $s>1$ represent the sub-Ohmic, Ohmic and super-Ohmic baths, respectively.
-This model is widely used to characterize the noise in nanomechanical devices Seoanez et al.
-(2007); Galland et al.
-(2008); Chen et al.
+This model is widely used to characterize the noise in nanomechanical devices Galland et al.
 (2008), superconducting circuits Magazzù et al.
-(2018); Thorwart et al.
 (2018); Kaur et al.
-(2021); Gulácsi and Burkard (2023), and proton-transfer reactions Cukier and Morillo (1989); Ohta et al.
+(2021), and proton-transfer reactions Ohta et al.
 (2006); Craig et al.
 (2007).
-The spin dynamics are expected to exhibit a phase transition from coherent oscillations to incoherent decay or localization as $s$ is decreased or $A$ is increased, which has been a subject of longstanding research using analytical tools Leggett et al.
+The spin dynamics are expected to exhibit a phase transition from coherent oscillations to incoherent decay or localization as $s$ is decreased or the amplitude $A$ is increased, which has been a subject of longstanding research using analytical tools Leggett et al.
 (1987); Bulla et al.
-(2003); Vojta et al.
 (2003) and classical simulations Wang and Thoss (2008); Duan et al.
 (2017); Strathearn et al.
-(2018); Kast and Ankerhold (2013); Nalbach and Thorwart (2013); Otterpohl et al.
-(2022); Nacke et al.
-(2023).
-We use a sum of Lorentzian lines from motional modes to approximate the spectral density $J_{\rm Legg}(\omega)$ Meier and Tannor (1999).
-To find the optimal set of parameters $\kappa$’s, $\Gamma$’s, and $\nu$’s in Eq.
-(3) that best represent $J_{\rm Legg}(\omega)$, well-established global optimization algorithms Wolz (2022) are used to minimize the cost function, defined as the integral of the absolute difference between $J_{\rm Legg}(\omega)$ and the sum of corresponding Lorentzian spectral densities over a given bandwidth sup .
-As a first step, we consider weak spin-bath coupling ($A=0.1$) and designed the bath spectrum for a limited bandwidth near resonance $\omega\in[0.9\Delta,1.1\Delta]$ using 2 Lorentzian peaks.
+(2018); Nalbach and Thorwart (2013).
+Recognizing that the spin exchanges energy with the bath only near its resonance, we use a sum of Lorentzian lines from motional modes to approximate the spectral density $J_{\rm Legg}(\omega)$ near the resonance frequency $\Delta$.
+We utilize well-established global optimization algorithms to find the optimal set of parameters ($\kappa_{l}$, $\Gamma_{l}$, and $\nu_{l}$ in Eq.
+(3)) that best represent $J_{\rm Legg}(\omega)$ sup .
+Using this approach, we consider weak spin-bath coupling ($A=0.1$) and match the spectral density within the target bandwidth $\omega\in[0.9\Delta,1.1\Delta]$ with 2 Lorentzian peaks.
 Figure 4a shows the obtained spectra from the modes (solid lines) that approximate sub-Ohmic ($s=0.5$), Ohmic ($s=1.0$), and super-Ohmic ($s=2.0$) spectral densities (dot-dashed lines), respectively.
 The results for the donor population dynamics interacting with a bath featuring these approximated spectral densities are shown in Fig.
-3b-d.
-We observe a crossover from near-coherent ($s=2$) to fully damped ($s=0.5$) oscillations in the simulated time scale depending on the level of noise in the bath.
-The experimental data match well with theoretical predictions, where the deviations in the $s=2$ case result from insufficient number of Trotterization steps sup .
-Figure 4: Simulating sub-Ohmic ($s=0.5$) bath with 10 oscillators.
-(a) Comparison between the sub-Ohmic spectral density (dot-dashed blue) and its approximation constructed by 10 noisy oscillators (solid red).
-Each dashed line represents the Lorentzian spectral density generated by each of these 10 oscillators.
-(b) Comparison of the population dynamics generated by the sub-Ohmic spectral density (dot-dashed blue) and its 10-mode approximation (solid red).
-Figure 3e compares the predicted population dynamics of the spin-boson model with ideal spectral density $J_{\rm Legg}(\omega)$ and the corresponding approximated spectral density.
-While the dynamics match very well for the Ohmic and super-Ohmic cases, the sub-Ohmic case shows a non-negligible deviation due to the gap in the two spectral densities within the bandwidth (insets in Fig.
-3a).
-In future experiments with a larger number of motional modes Porras et al.
-(2008), we expect to match a wider range of spectral densities and thus more accurately simulate the sub-Ohmic spin-boson model, as suggested by the 10-mode numerical example in Fig.
-4.
+4b-d.
+We observe a crossover from near-coherent ($s=2$) to fully damped ($s=0.5$) oscillations in the simulated time scale, owing to the different levels of spectral density $J(\omega=\Delta)$ near resonance.
+The observed behavior qualitatively agrees with the predictions of the Leggett model: for super-Ohmic case, coherence is maintained due to low spectral density near resonance, while for sub-Ohmic case, large spectral density induces strong damping to the point where the system is localized.
+We note that the deviations in the $s=2$ case in Fig.
+4d result from insufficient number of Trotterization steps sup .
 Vibration-assisted energy transfer
@@
 Here we set the spin parameters as $\Delta=30\>\text{cm}^{-1}$ and $\epsilon=100\>\text{cm}^{-1}$.
-In the absence of coupling to a vibrational mode ($\kappa=0\>\text{cm}^{-1}$), energy transfer between the detuned spin states is suppressed; however, for nonzero $\kappa$, the energy transfer can be activated if the mode frequency $\nu$ meets the resonance condition.
-Figure 5: Simulations of vibrationally-assisted energy transfer (VAET).
+In the absence of coupling to a vibrational mode ($\kappa=0\>\text{cm}^{-1}$), energy transfer between the detuned spin states is suppressed; however, for nonzero $\kappa$, the energy transfer can be activated if the mode frequency $\nu$ meets the resonance condition (single mode’s index $l=1$ omitted).
+Figure 5a shows the experimental data for simulating the coherent spin-single oscillator model $\hat{H}_{D}$ with a relatively large coupling strength ($\kappa=30\>\text{cm}^{-1}$).
+When $\nu$ satisfies the resonance condition of VAET ($\nu=\sqrt{\Delta^{2}+\epsilon^{2}}/k\approx 104\>\text{cm}^{-1}/k$ where $k$ is an integer), energy transfer occurs Gorman et al.
+(2018).
+Otherwise, energy remains localized at the donor state.
+Figure 5b shows the impact of initial temperature of the oscillator on the resonant VAET ($\nu=104\>\text{cm}^{-1}$).
+We create baths of two temperatures 0 K and 300 K corresponding to initial average phonon number $\bar{n}=0$ and $2.0$, respectively sup .
+The energy transfer is faster for higher $\bar{n}$ in the beginning as more vibrational quanta assist the energy transfer, but it also leads to a rapid decay of oscillations over time as the coherence is lost.
+In Fig.
+5c, we simulate the dephased spin-oscillator model with $\Gamma=10\>\text{cm}^{-1}$ and various values of $\kappa$, to study the impact of dissipative coupling.
+We note that in the weak-$\kappa$ regime (or short evolution time), the results agree with the theoretical predictions of the damped spin-oscillator model, which is equivalent to those of the spin-boson model with $J(\omega)=J_{\rm Lo}(\omega)$ Lemmer et al.
+(2018); Tamascelli et al.
+(2018); however, as $\kappa$ is increased, the dephased and damped models’ dynamics deviate.
+This is expected because unlike the damped model, the dephased model does not extract energy from the spin-oscillator system, so the donor population will not decay towards zero even after a long time evolution when coupled to a single-oscillator bath.
+This example demonstrates that our method provides the ability to simulate a wider range of dissipative channels in the study of open system dynamics.
+Figure 5: Simulations of vibration-assisted energy transfer (VAET).
 We set $\epsilon=100\>\text{cm}^{-1}$ and $\Delta=30\>\text{cm}^{-1}$.
 (a) (b) Time evolution of the donor-state population for coherent VAET models with $\kappa=30\>\text{cm}^{-1}$.
-Solid (dashed) curves depict theoretical predictions of the target (noise-aware) model.
+Solid and dashed curves depict theoretical predictions of the ideal target and the noise-aware model, respectively.
 Circles represent experimental data and error bars are standard deviation over 100 repetitions.
@@
 In (b), $\nu$ is fixed to the resonant value $104\>\text{cm}^{-1}$ and two temperatures 0 K and 300 K, which correspond to $\bar{n}=0$ and 2.0, respectively, are simulated.
-(c) Time evolution of the donor-state population for noisy VAET models with $\nu=104\>\text{cm}^{-1}$ and $\Gamma=10\>\text{cm}^{-1}$ at 0 K.
-Solid (dashed) curves depict theoretical predictions of the spin-boson (dephased spin-oscillator) model.
+(c) Time evolution of the donor-state population for dissipative VAET models with $\nu=104\>\text{cm}^{-1}$ and $\Gamma=10\>\text{cm}^{-1}$ at 0 K.
+Solid and dot-dashed curves depict theoretical predictions of the dephased spin-oscillator model and the damped spin-oscillator model, respectively.
 Circles denote experimental data.
@@
 For $\kappa=10$ and $20\>\text{cm}^{-1}$, error bars indicate the standard deviation over 20 trials, each trial performed with randomly drawn set of parameters and repeated 100 times.
-Figure 5a shows the experimental data for simulating the coherent spin-oscillator model $\hat{H}_{SO}$ with a relatively large coupling strength ($\kappa=30\>\text{cm}^{-1}$).
-When $\nu$ satisfies the resonance condition of VAET ($\nu=\sqrt{\Delta^{2}+\epsilon^{2}}/k\approx 104\>\text{cm}^{-1}/k$ where $k$ is an integer), energy transfer occurs Gorman et al.
-(2018).
-Otherwise, energy remains localized at the donor site.
-Figure 5b, shows the impact of initial temperature of the oscillators on the resonant VAET ($\nu=104\>\text{cm}^{-1}$).
-We simulate baths of two temperatures 0 K and 300 K corresponding to initial average phonon number $\bar{n}=0$ and $2.0$, respectively (see Methods).
-The energy transfer is faster for higher temperature bath in the beginning, as more vibrational quanta assist the energy transfer.
-However, high temperature leads to a rapid decay of oscillations over time as the coherence is lost.
-In Fig.
-5c, we simulate the dephased spin-oscillator model with $\Gamma=10\>\text{cm}^{-1}$ and various values of $\kappa$, to study the impact of dissipative coupling.
-We note that as $\kappa$ is increased, the results deviate more from the theoretical predictions of the damped spin-oscillator model, which is non-perturbatively equivalent to those of the spin-boson model with $J(\omega)=J_{\rm Lo}(\omega)$ Lemmer et al.
-(2018); Tamascelli et al.
-(2018).
-This is consistent with the expectation that the method of using the dephased spin-oscillator model to approximate the spin-boson model is valid only in the weak-$\kappa$ regime.
-While the short-time dynamics between the two models match as the 2-point correlation functions are the same, the long-time dynamics deviate due to mismatch in the higher-order correlation functions for higher coupling strengths sup .
 Discussions and outlook
-Simulations of large-scale spin-boson models with strong coupling in the presence of dissipation is crucial yet challenging problem in physical Caldeira and Leggett (1983); Weiss et al.
-(1987); Schindler et al.
-(2022), chemical Garg et al.
-(1985); Gilmore and McKenzie (2005), and biological Huelga and Plenio (2013); Jang and Mennucci (2018) systems.
-While classical simulation approaches continue to improve with advancements in numerical methods, they encounter fundamental challenges when dealing with large quantum systems that span exponentially large Hilbert space MacDonell et al.
+While classical simulation approaches to open dissipative systems continue to improve with advancements in numerical methods, they encounter fundamental challenges when dealing with large quantum systems that span exponentially large Hilbert space MacDonell et al.
 (2021); Kang et al.
@@
 (2024).
-The current state-of-the-art trapped ion systems offer several dozens of highly coherent spins and bosonic modes that are very densely coupled with each other, the complete quantum description of which is totally intractable with classical computers of today Cetina et al.
+The current state-of-the-art trapped-ion systems offer several dozens of highly coherent spins and bosonic modes that are very densely coupled with each other, the complete quantum description of which is totally intractable with classical computers of today Cetina et al.
 (2022); Chen et al.
-(2023b).
-Developing various control methods over these quantum systems to study important physical, chemical, biological and material systems can open up a new computational framework in science.
-The methods developed in this work utilizes fully programmable control methods to simulate dissipative processes.
-This approach can readily be applied to existing simulations of chemical processes to include coupling to environment Whitlow et al.
+(2023b); DeCross et al.
+(2024).
+Developing various control methods over these quantum systems to study important physical, chemical, biological and material systems of critical interest will open up a new computational framework in science.
+The methods developed in this work utilize fully programmable control methods to simulate dissipative processes.
+The dephased oscillator method can be used to realize non-Gaussian bath models Barthel and Zhang (2022), relevant for studying physical processes such as phase transitions in strongly-coupled spin-boson models Leggett et al.
+(1987); Wang and Thoss (2008); Duan et al.
+(2017); Strathearn et al.
+(2018); Nalbach and Thorwart (2013) and role of bosonic dephasing in molecular energy transfer Stock and Domcke (1990); Banin et al.
+(1994).
+The capability to measure the probability distribution of the motional modes Flühmann and Home (2020); Jia et al.
 (2022); Valahu et al.
-(2023); Sun et al.
-(2023); Zhang et al.
-(2024); Li et al.
-(2021).
-The dephased oscillator method can be used to realize non-Gaussian bath models Linowski et al.
-(2022); Barthel and Zhang (2022), relevant for studying physical processes such as phase transitions in strongly-coupled spin-boson models Duan et al.
-(2017); Otterpohl et al.
-(2022) and role of bosonic dephasing in molecular energy transfer Grad et al.
-(1988); Goldsmith et al.
-(2006); Plenio and Huelga (2008); Hayes et al.
-(2010); Moreira et al.
-(2020); Yoneda et al.
-(2021).
+(2023); Whitlow et al.
+(2023) may further reveal the effects of intricate vibronic interactions in biological and chemical reaction dynamics.
 High accuracy simulations will require development of efficient characterization methods Kang et al.
-(2023b) and novel control protocols in the trapped ion system sup ; Lobser et al.
+(2023b) and novel control protocols in the trapped-ion system sup ; Lobser et al.
 (2023) to implement various interaction terms in the target Hamiltonian.
@@
 (2018); Schlawin et al.
-(2021), the long-chain trapped ion systems can provide a versatile platform for simulating dynamics of complex open quantum systems.
+(2021), the long-chain trapped-ion systems can provide a versatile platform for simulating dynamics of complex open quantum systems.
 During the preparation of our manuscript we became aware of a complementary work So et al.
 (2024) which simulates the electron-transfer model using sympathetic cooling on the trapped ions’ motional mode.
-Acknowledgements.
-The authors thank Yichao Yu for his valuable discussions about the phase tracking method, Daniel Lobser for his valuable suggestion in controlling the RFSoC control software, and Thomas Barthel for his course on open quantum systems that inspired this work.
-This research is funded by the Office of the Director of National Intelligence - Intelligence Advanced Research Projects Activity, through the ARO contract W911NF-16-1-0082 (provided the experimental apparatus used in the demonstration), the DOE BES award DE-SC0019400 (theoretical modeling), and the NSF Quantum Leap Challenge Institute for Robust Quantum Simulation Grant No.
+Acknowledgments: The authors thank Yichao Yu for his valuable discussions about the phase tracking method, Daniel Lobser for his valuable suggestion in controlling the RFSoC control software, and Thomas Barthel for his course on open quantum systems that inspired this work.
+Funding: This research is funded by the Office of the Director of National Intelligence - Intelligence Advanced Research Projects Activity, through the ARO contract W911NF-16-1-0082 (provided the experimental apparatus used in the demonstration), the DOE BES award DE-SC0019400 (theoretical modeling), and the NSF Quantum Leap Challenge Institute for Robust Quantum Simulation Grant No.
 OMA-2120757 (experimental concepts and validation).
@@
 Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator (data analysis and validation).
-Methods
-Experimental implementation
-Figure 6: Experimental setup and instruction of operations.
-(a) Schematic of the experimental setup illustrates the orientation of relevant fields.
-The magnetic field and the direction of the individual beam is perpendicular to the trap surface, while the global beam propagates is parallel to it.
-The differential wave vector $\Delta\textbf{k}$ is therefore oriented at a 45-degree angle to the trap surface, aligning with one of the radial axes.
-The global beam comprises two tones driving the blue (BSB) and red (RSB) sideband transitions, respectively.
-(b) Simplified energy levels of the 171Yb+ ion and the laser pulses used for implementing the on-resonance spin-dependent kick (SDK) operation.
-(c) Visualization of the motional-state evolution in phase space when a SDK operation of duration $\tau$ is applied, where the qubit is in the eigenstate of the operator $\hat{\sigma}_{\phi_{s}}\equiv\hat{\sigma}_{+}e^{i\phi_{s}}+\hat{\sigma}_{-}e^{-i\phi_{s}}$.
-Here, $\hat{\sigma}_{+}$ ($\hat{\sigma}_{-}$) is the raising (lowering) operator of the qubit, $\phi_{s}$ ($\phi_{m}$) is the spin (motion) phase of the laser beams, $\tilde{\Omega}$ is the Rabi frequency of the sideband transition, and $\delta_{m}$ is the motion detuning.
-(d) Instruction of operations for simulating the dephased spin-oscillator model, which approximates the spin-boson model.
-‘H’ represents the Hadamard gate on the qubit.
-‘Kick (heat)’ represents the SDK operations with random $\phi_{m}$ that prepare the thermal state on average.
-‘SQ’ (‘Kick (coh/dep)’) represents the single-qubit rotations (SDK operations) that simulate the first (second) term in Eq.
-(5).
-Coherent (dephased) spin-oscillator model is simulated by using a fixed (randomly drawn) $\delta_{m}$ value at each step.
-The simulations of spin-boson models are performed using a linear chain of 171Yb+ ions confined in a micro-fabricated surface trap Revelle (2020), as illustrated in Fig.
-6a.
-The spin states are encoded in the hyperfine energy levels in the 2S1/2 manifold (qubit states) as |$\ket{0}\equiv\ket{F=1;m_{F}=0}$ and |$\ket{1}\equiv\ket{F=0;m_{F}=0}$.
-The bosonic bath modes are encoded as the collective radial motional modes of the ion chain.
-Details of our experimental setup can be found in Ref.
-Wang et al.
-(2020).
-Through the sequential application of Doppler cooling, EIT cooling, and sideband cooling techniques, the zigzag motional mode of the ion chain can be cooled to $\bar{n}=0.036(16)$.
-Following this, a $370$ nm laser is utilized to pump the ions to their ground state |$\ket{1}$.
-Transitions between the qubit states are performed using stimulated Raman transitions, driven by a pair of laser beams Wang et al.
-(2020).
-Acousto-optic modulators are employed to manipulate the frequency and phase of each laser beam, thereby enabling single-qubit gates: we apply a bit flip ($X$) gate to initialize the ion to |$\ket{0}$ state.
-The spin-oscillator coupling is simulated using the spin-dependent kick (SDK) operation induced by simultaneous blue- and red-sideband transitions as shown in Fig.
-6b.
-Figure 6c visualizes how the motional state evolves in phase space when a SDK operation of duration $\tau$, spin (motion) phase $\phi_{s}$ ($\phi_{m}$), motion detuning $\delta_{m}$, and sideband Rabi frequency $\tilde{\Omega}$ is applied Lee et al.
-(2005).
-Subsequent to the final operation, the qubit-state population is accessed via the state-dependent detection technique.
-We strategically use the center ion of a 7-ion chain, as it exhibits negligible coupling to the 2nd, 4th, and 6th modes (See Fig.
-7).
-This arrangement maximizes the effective spacing between the motional-mode frequencies and therefore suppresses cross-mode coupling (driving modes that are not targeted by the SDK operation).
-Figure 7: Ion-Mode coupling strengths of a 7-ion chain. Motional-mode frequencies are shown in the lower right corner of each panel.
-Trotterization
-The evolution of $\hat{H}_{SO}$ in Eq.
-(2) is mapped to a sequence of single-qubit rotations and SDK operations via Trotterization.
-After conjugating $\hat{H}_{SO}$ with a Hadamard gate on the qubit, the Hamiltonian can be described in the interaction picture as
-$\hat{H}_{I}(t)=\frac{\epsilon}{2}\hat{\sigma}_{\Delta}(t)+\frac{\kappa}{2}\hat{\sigma}_{\Delta}(t)(\hat{b}e^{-i\nu t}+\hat{b}^{\dagger}e^{i\nu t}),$ (5)
-where $\hat{\sigma}_{\Delta}(t)\equiv\hat{\sigma}_{+}e^{i\Delta t}+\hat{\sigma}_{-}e^{-i\Delta t}$ and $\hat{\sigma}_{+}$ ($\hat{\sigma}_{-}$) is the raising (lowering) operator of the qubit.
-The evolution with respect to $\hat{H}_{I}(t)$ is discretized into $N$ time steps, each consisting of a single-qubit rotation and a SDK operation that simulate the first and second term, respectively.
-Specifically, for the SDK operation at the $j$-th time step, the spin (motion) phase of the laser beams Lee et al.
-(2005) is set as $\Delta t_{j}$ ($\nu t_{j}$), where $t_{j}\equiv(j-\frac{1}{2})T/N\tau$ and $\tau$ is the duration of the operation for each time step.
-This is equivalent to setting the symmetric (anti-symmetric) detuning from the sideband transitions, referred to as the spin (motion) detuning, as $\Delta T/N\tau$ ($\nu T/N\tau$) at all time steps.
-This method can be straightforwardly extended to the case where multiple modes are used; however, the laser phase needs to be tuned to track the phase of each motional mode correctly sup .
-This approach can be readily extended to simulating more complicated molecular models consisting of many electronic states and vibrational modes Kang et al.
-(2024).
-Randomized operations
-Randomizing the control parameters is a key ingredient in our approach to simulating dissipation mechanisms in spin-boson models.
-By taking an average of many coherent evolutions each involving a stochastically varying parameter, we realize both heating and dephasing as described by the Lindblad master equation Cai and Barthel (2013); Chenu et al.
-(2017).
-Heating (dephasing) is used to tune the temperature (spectral linewidth) of the bath.
-First, the bosonic modes at finite temperature is created by modeling heating process where the modes interact with an infinite-temperature external bath for a finite duration, described by aa pair of Lindblad operators $\hat{L}_{1}=\sqrt{\Gamma^{\prime}}\hat{b}$ and $\hat{L}_{2}=\sqrt{\Gamma^{\prime}}\hat{b}^{\dagger}$.
-We apply $N^{\prime}$ resonant ($\delta_{m}=0$) SDK operations of duration $\tau^{\prime}=4\Gamma^{\prime}/\tilde{\Omega}^{2}$, spin phase $\phi_{s}=0$, and motion phase $\phi_{m}$ randomly drawn from $[0,2\pi)$ for each operation.
-These stochastic operations are applied on the initial qubit-mode composite state |$\ket{+,n=0}$; note that $\ket{+}$ is an eigenstate of the spin operator with $\phi_{s}=0$, and thus the qubit is decoupled from these operations.
-When averaged over many trials of coherent evolutions, each executed with a distinct set of $N^{\prime}$ random $\phi_{m}$ values, this procedure prepares the thermal state with an average phonon number $\bar{n}=N^{\prime}\Gamma^{\prime}\tau^{\prime}$.
-This can be intuitively understood as the thermal state represented as an ensemble of randomly displaced coherent states (derivation provided in Supplementary Material sup ).
-For the experimental data in Fig.
-1a and Fig.
-5b, the average phonon number $\bar{n}$ of the initial thermal state is tuned by varying the duration $\tau^{\prime}$ of SDK operations, where the number of steps $N^{\prime}$ is fixed to 5.
-The errors in $\bar{n}$ due to finite number of steps is discussed in the Supplementary Material sup .
-Next, dephasing during the time evolution, which is described by a single Lindblad operator $\hat{L}_{1}=\sqrt{\Gamma}\hat{b}$, is induced by giving random offsets to the motion detuning $\delta_{m}$ of the $N$ Trotterized SDK operations.
-Specifically, an offset to $\delta_{m}$ is assigned for each of the $N$ steps, with values drawn from a normal distribution characterized by a mean of zero and a standard deviation of $\sqrt{\Gamma T/N}/\tau$ sup .
-The qubit population dynamics averaged over many trials of coherent evolutions, each performed with a set of $N$ random $\delta_{m}$ values, give the dynamics with respect to the dephased spin-oscillator model.
-The entire instruction of operations is shown in Fig.
-6d.
-In practice, the dephasing strength $\Gamma$ is varied by tuning the standard deviation of the random offsets to $\delta_{m}$ at each Trotterization step.
-This enables tuning the FWHM of each Lorentzian peak in the bath’s spectral density.
-The errors in the simulated dephasing strength due to finite number of Trotterization steps is discussed in the Supplementary Material sup .
+Author contributions: K.S., M.K., K.R.B.
+and J.K.
+conceived the idea and developed detailed methodology for experiments and modeling, K.S.
+and G.S.
+performed the experiments, M.K., H.N.
+and D.N.B performed modeling and simulations, K.S., M.K., K.R.B.
+and J.K.
+analyzed and validated the data, K.R.B.
+and J.K.
+supervised the project, K.S.
+and M.K.
+wrote the original draft, and K.S., M.K.
+and J.K.
+revised the manuscript.
+All authors discussed the results and reviewed the final manuscript.
+Competing Interests: J.K.
+and K.R.B are scientific advisors to IonQ, Inc., and J.K.
+is a shareholder of IonQ, Inc.
+Data and materials availability: All data needed to evaluate the conclusions in the paper are present in the paper and/or Supplementary Materials.
+We will provide raw data to anyone who might be interested.